### v2

# Rankify: A Comprehensive Python Toolkit for Retrieval, Re-Ranking, and Retrieval-Augmented Generation

Abstract.
Retrieval, re-ranking, and retrieval-augmented generation (RAG) are critical components of modern natural language processing (NLP) applications in information retrieval, question answering, and knowledge-based text generation. However, existing solutions are often fragmented, lacking a unified framework that easily integrates these essential processes. The absence of a standardized implementation, coupled with the complexity of retrieval and re-ranking workflows, makes it challenging for researchers to compare and evaluate different approaches in a consistent environment. While existing toolkits such as Rerankers and RankLLM provide general-purpose reranking pipelines, they often lack the flexibility required for fine-grained experimentation and benchmarking. In response to these challenges, we introduce Rankify, a powerful and modular open-source toolkit designed to unify retrieval, re-ranking, and RAG within a cohesive framework. Rankify supports a wide range of retrieval techniques, including dense and sparse retrievers, while incorporating state-of-the-art re-ranking models to enhance retrieval quality. Additionally, Rankify includes a collection of pre-retrieved datasets to facilitate benchmarking, available at Huggingface [^1] https://huggingface.co/datasets/abdoelsayed/reranking-datasets-light. To encourage adoption and ease of integration, we provide comprehensive documentation [^2] http://rankify.readthedocs.io/, an open-source implementation on GitHub [^3] https://github.com/DataScienceUIBK/rankify, and a PyPI package for effortless installation [^4] https://pypi.org/project/rankify/. By providing a unified and lightweight framework, Rankify allows researchers and practitioners to advance retrieval and re-ranking methodologies while ensuring consistency, scalability, and ease of use.
Neural IR, Dense Retrieval, Sparse Retrieval, Re-ranking, Toolkit
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Information systems Evaluation of retrieval resultsccs: Information systems Presentation of retrieval resultsccs: Information systems Top-k retrieval in databasesccs: Information systems Rank aggregation
Figure 1. Rankify logo.
1. Introduction
Information retrieval (IR) systems (Singhal et al., 2001; Chowdhury, 2010) are fundamental to many natural language processing (NLP) applications (Wang et al., 2024; Liu et al., 2024), including question-answering (Khamnuansin et al., 2024; Abdallah et al., 2024b), search engines (Croft et al., 2010; Xiong et al., 2024), and knowledge-based generation (Long et al., 2024). These systems often rely on a two-stage pipeline: a retriever that efficiently identifies a set of candidate documents and a re-ranker that refines these results to maximize relevance to the query (Yates et al., 2021; Abdallah et al., 2024a; Nogueira et al., 2019; Ren et al., 2021). This approach has proven highly effective, with retrieval and re-ranking methods achieving state-of-the-art performance across diverse NLP benchmarks.
Retrievers form the backbone of information retrieval systems by identifying a subset of documents relevant to a user’s query. The retrieval landscape includes sparse, dense, and hybrid methods. Sparse retrievers, such as BM25 (Robertson et al., 1995), rely on exact term matching by representing queries and documents as high-dimensional sparse vectors. These methods are highly effective when query terms closely align with document content but struggle to capture semantic relationships. Dense retrievers, including models like DPR (Yates et al., 2021), BPR (Yamada et al., 2021), ColBERT (Santhanam et al., 2021), BGE (Chen et al., 2023), Contriever (Izacard et al., 2021) and ANCE (Xiong et al., 2021a), overcome this limitation by encoding queries and documents into low-dimensional dense vectors using neural networks, enabling retrieval based on semantic similarity even when lexical overlap is minimal. However, dense retrieval requires significant computational resources for training and inference. Hybrid retrieval methods (Yang et al., 2019; Chen et al., 2022) integrate the strengths of both approaches, combining the precision of sparse retrievers with the semantic understanding of dense models to achieve a more balanced and robust retrieval system.
Toolkit # Datasets # Retrievers # Re-rankers # RAG Methods Modular Design Automatic Evaluation Corpus
FlashRAG (Jin et al., 2024) 32 3 2 12
FastRAG (Abane et al., 2024) 0 1 3 7
AutoRAG (Kim et al., 2024a) 4 2 7 1
LocalRQA (Yu et al., 2024) 0 2 0 0
Rerankers (Clavié, 2024) 0 0 15 0
Rankify 40 7 24 3+
Table 1. Comparison with other Retriever, Re-ranking, and RAG toolkits. Modular Component refers to whether the toolkit is composed of modular components. Automatic Evaluation indicates if the toolkit provides automated evaluation capabilities to assess the performance of various datasets. Corpus Helper shows if the toolkit offers auxiliary tools for processing corpus, including cleaning and chunking.
Re-ranking techniques enhance the initial retrieval results by ensuring the most relevant documents appear at the top. These methods are categorized into three main approaches: (1) Pointwise reranking (Nogueira et al., 2019; Abdallah et al., 2025b, a; Sachan et al., 2022) treats reranking as a regression or classification task, assigning independent relevance scores to each document. (2) Pairwise reranking (Huang and Chen, 2024; Qin et al., 2023; Sinhababu et al., 2024) refines ranking by comparing document pairs and optimizing their relative order based on relevance. However, this approach treats all document pairs equally, sometimes improving lower-ranked results at the expense of top-ranked ones. (3) Listwise reranking (Sun et al., 2023a; Pradeep et al., 2023b; Yoon et al., 2024) considers the entire document list, prompting LLMs with the query and a subset of candidates for reranking. Due to input length limitations, these methods tend to employ a sliding window strategy, progressively refining rankings from back to front.
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Mozafari et al., 2024; Abdallah and Jatowt, 2023; Gao et al., 2023) enhances language models by integrating retrieval and generation, making them more effective in knowledge-intensive tasks. Rather than relying solely on pre-trained knowledge, RAG dynamically retrieves relevant documents from external sources during inference, incorporating them into the generation process.
The growing complexity and diversity of retrieval, re-ranking, and RAG methods pose significant challenges for benchmarking, reproducibility, and integration. Existing toolkits, such as Pyserini (Lin et al., 2021), Rerankers (Clavié, 2024) and RankLLM (Pradeep et al., 2023a), often lack flexibility, enforce rigid implementations, and require extensive preprocessing, making them less suitable for research-driven experimentation. Additionally, retrieval and re-ranking datasets are scattered across different sources, complicating evaluation and comparison. To address these challenges, we introduce Rankify, an open-source framework that unifies retrieval, re-ranking, and RAG into a modular and extensible ecosystem with its logo shown in Figure 1. Rankify supports diverse retrieval techniques, integrates state-of-the-art re-ranking models, and provides curated pre-retrieved datasets to streamline experimentation. Designed with flexibility, it enables researchers to efficiently build, evaluate, and extend retrieval pipelines while ensuring consistency in benchmarking.
The main contributions of this paper are as follows:
  • Curated retrieval datasets and precomputed embeddings: Rankify provides 40 datasets with 1,000 pre-retrieved documents per query across various domains (QA, dialogue, entity linking, etc.). It also includes pre-computed Wikipedia and MS MARCO corpora for multiple retrievers, eliminating preprocessing overhead.
  • Diverse retriever and re-ranking support: Rankify integrates dense (DPR, ANCE, BPR, ColBERT, BGE, Contriever) and sparse (BM25) retrievers, along with 24 state-of-the-art re-ranking models, enabling seamless retrieval and ranking experimentation.
  • RAG integration and evaluation tools: Rankify bridges retrieval, re-ranking, and RAG by passing retrieved documents to LLMs for evaluation.
  • Comprehensive evaluation tools: Rankify offers a diverse range of evaluation metrics and tools for retrieval and question answering. The framework is accompanied by extensive online documentation, making it easy for users to explore its features. Additionally, Rankify is freely available on PyPI and GitHub, ensuring accessibility for researchers and practitioners.
2. Related Work
2.1. Retrievers
Retrieval models have evolved significantly, transitioning from traditional sparse retrievers to sophisticated dense and hybrid retrieval techniques. Sparse retrievers, such as BM25 (Robertson et al., 1995), remain widely used due to their strong lexical matching capabilities and generalization. However, they struggle with capturing semantic relationships, leading to the rise of dense retrieval methods (Lee et al., 2019), which leverage pre-trained neural encoders to generate embeddings for queries and documents. Notable advancements, such as DPR (Yates et al., 2021) and ANCE (Xiong et al., 2021a), improve retrieval effectiveness by employing contrastive training with in-batch negatives (Gillick et al., 2019) and hard-negative mining (Karpukhin et al., 2020; Xiong et al., 2021b). Multi-vector approaches like ColBERT (Khattab and Zaharia, 2020) enable fine-grained interaction between queries and documents, albeit at the cost of increased computational overhead. To address the limitations of dense retrievers, hybrid models (Gao et al., 2021; Ma et al., 2022) combine sparse and dense approaches, utilizing both term matching and semantic understanding for enhanced retrieval performance.
Recent research has also explored retrieval enhancements through knowledge distillation, curriculum learning, and pre-training. Distillation techniques transfer knowledge from cross-encoders or teacher retrievers to student models, improving dense retriever effectiveness (Qu et al., 2021; Chen et al., 2021). Curriculum learning strategies, which progressively train retrievers from easy to hard samples (Zeng et al., 2022; Lin et al., 2022a), have been shown to enhance both supervised and zero-shot retrieval effectiveness. Meanwhile, pre-training methods such as contrastive learning (Izacard et al., 2021) and masked autoencoding (Gao and Callan, 2021; Liu and Shao, 2022) have significantly improved the generalization of retrieval models. Additionally, approaches like query augmentation (Mao et al., 2021; Ma et al., 2021) generate synthetic training data to enhance retrieval robustness.
2.2. Re-Ranking Methods
Re-ranking plays a crucial role in refining retrieved document lists to improve ranking quality. Traditional approaches include pointwise learning-to-rank (LTR) methods, where each document is scored independently (Zhu et al., 2021; Cinar and Renders, 2020), and pairwise methods, which optimize the relative ordering of document pairs (Cao et al., 2007; Jia et al., 2021). Listwise approaches, such as LambdaRank and ListNet, directly optimize ranking metrics like Discounted Cumulative Gain (DCG) to enhance top-ranked results (Burges, 2010; Liu et al., 2017; Cao et al., 2007). More recently, deep neural networks and large language models (LLMs) have revolutionized re-ranking, enabling context-aware ranking adjustments. Cross-encoder models, which jointly encode queries and candidate documents, have demonstrated significant improvements in ranking accuracy by capturing fine-grained interactions between them. Additionally, zero-shot re-ranking approaches leverage models like GPT-4 for listwise re-ranking without task-specific training (Achiam et al., 2023; Sun et al., 2023c), while knowledge distillation techniques such as RankZephyr fine-tune smaller models to approximate LLM performance with reduced computational cost (Pradeep et al., 2023b).
Re-ranking is also increasingly integrated into retrieval-augmented generation (RAG) systems, where refining retrieved documents before response generation ensures higher relevance and coherence in generated outputs. These advancements highlight the transition from traditional ranking techniques to sophisticated, model-driven re-ranking approaches that enhance retrieval quality in diverse NLP applications. Recent innovations in re-ranking include leveraging large language models (LLMs) such as GPT-4 (Achiam et al., 2023) for zero-shot and listwise re-ranking (Sun et al., 2023c). Other approaches, such as RankZephyr (Pradeep et al., 2023b), utilize knowledge distillation to fine-tune smaller models based on LLM outputs, enabling efficient and effective re-ranking.
2.3. Retrieval-Augmented Generation (RAG)
Retrieval-Augmented Generation (RAG) combines retrieval and generation to enhance language models in knowledge-intensive tasks. Early RAG systems (Lewis et al., 2020) retrieve context documents relevant to a query and use these as input for generation, enabling models to answer queries beyond their training data. More recent approaches explore the integration of retrieval with both encoder-decoder (Izacard and Grave, 2020) and decoder-only (Raffel et al., 2020) architectures, optimizing performance for a variety of downstream tasks.
One key area of focus in RAG is mitigating hallucinations and improving the factuality of generated responses (Manakul et al., 2023). Techniques such as self-consistency (Wang et al., 2022b) and noisy context filtering (Fang et al., 2024) have been proposed to enhance the reliability of RAG outputs. However, the quality of retrieved documents remains a critical factor in the overall effectiveness of RAG pipelines. Recent works emphasize the importance of combining retrieval with robust re-ranking techniques to ensure that the retrieved context is both relevant and precise (Kamalloo et al., 2023).
2.4. Frameworks
Several frameworks have been developed to support research and development in retrieval and re-ranking systems. rerankers (Clavié, 2024) is a lightweight Python library that provides a unified interface to commonly used re-ranking approaches, simplifying the integration of various methods into retrieval pipelines. Similarly, RankLLM [^5] https://github.com/castorini/rank_llm/tree/main offers a Python toolkit focused on reproducible information retrieval research, emphasizing listwise re-ranking with large language models. Additionally, FlashRAG (Jin et al., 2024) and AutoRAG (Kim et al., 2024b) introduce a modular toolkit designed for efficient retrieval-augmented generation (RAG) research, aiming to enhance flexibility and extensibility in experimental settings. These frameworks, along with others like LangChain (Chase, 2022), LlamaIndex (Liu, 2022), and DSPy (Khattab et al., 2023), provide researchers and practitioners with versatile tools to develop and evaluate retrieval and RAG applications. A comparison of retrieval, re-ranking, and RAG toolkits is presented in Table 1, highlighting the advantages of Rankify in dataset diversity, retriever and re-ranker support, and modularity.
Figure 2. The architecture of the Rankify, showing the interplay between its core modules: Datasets, Retrievers, Re-Rankers, and RAG Evaluation. Each module operates independently while seamlessly integrating with others, enabling end-to-end retrieval and ranking workflows.
3. Rankify
Modern information retrieval systems rely on a combination of retrieval, ranking, and generative techniques to surface relevant content and generate accurate responses. However, existing solutions are often fragmented, requiring researchers and practitioners to integrate multiple toolkits to experiment with different retrieval and ranking strategies. Rankify addresses this challenge by offering an end-to-end retrieval and ranking ecosystem, enabling seamless experimentation across a wide range of retrieval, re-ranking, and retrieval-augmented generation (RAG) models.
Unlike existing retrieval and ranking frameworks that focus on a single aspect of the pipeline, Rankify is designed to be modular, extensible, and lightweight, allowing users to plug in different retrievers, ranking models, and RAG techniques with minimal effort. Built in Python, it leverages machine learning libraries such as TensorFlow (Abadi et al., 2016), PyTorch (Paszke et al., 2019), Spacy (Honnibal et al., 2020), and Pyserini (Lin et al., 2021) . The framework is available on PyPI [^6] https://pypi.org/project/rankify/, making installation straightforward:
\$ pip install rankify
To ensure accessibility and reproducibility, Rankify supports both precomputed and real-time retrieval workflows. Users can work with pre-retrieved datasets for benchmarking or perform live document retrieval and re-ranking on large-scale corpora. The framework includes ready-to-use indices for Wikipedia and MS MARCO, eliminating the need for time-consuming indexing. Additionally, it supports different retrievers, giving users the flexibility to test different retrieval paradigms without modifying their pipeline.
Rankify consists of four core modules: Datasets, Retrievers, Re-Rankers, and RAG. The Datasets module provides standardized access to benchmark datasets, including Natural Questions, TriviaQA, and HotpotQA, as well as the ability to create custom datasets. The Retrievers module integrates state-of-the-art retrieval methods, including BM25, DPR, ANCE, BGE, and ColBERT, allowing users to experiment with different retrieval strategies. The Re-Rankers module includes 24 ranking models with 41 sub-methods, supporting pointwise, pairwise, and listwise re-ranking techniques. Finally, the RAG module enables retrieval-augmented generation using models such as LLAMA, GPT, and T5, supporting in-context learning, Fusion-in-Decoder (FiD), and retrieval-augmented latent memory (RALM).
A key advantage of Rankify is its unified API, which abstracts away implementation details, allowing users to experiment with different models and configurations using a simple interface. Figure 2 illustrates the architecture of Rankify, demonstrating how retrieval, ranking, and generation components interact within the pipeline. The framework is designed to be scalable and adaptable, making it suitable for both academic research and real-world applications in search, question answering, and document ranking.
To encourage collaboration and community contributions, Rankify is open-source and actively maintained [^7] https://github.com/abdoelsayed2016/reranking. Extensive documentation [^8] http://rankify.readthedocs.io/ and inline docstrings provide clear guidance on using the framework, ensuring ease of adoption for both beginners and experienced researchers. The next sections explore its capabilities in greater detail, covering retrieval benchmarks, ranking evaluations, and retrieval-augmented generation results.
Figure 3. Summary of datasets used in our Rankify. The figure illustrates the dataset sizes across different categories. The bubble size represents the total number of samples (Train + Dev + Test) for each dataset. Categories include QA, Multi-Hop QA, Long-Form QA, Multiple-Choice, Entity-Linking, Slot Filling, Fact Verification, Dialog Generation, Summarization, and Other specialized datasets. The log scale is used for better visualization of dataset size variations.
Figure 4. Schema of the Dataset class, illustrating the objects used to represent a dataset in Rankify.
3.1. Prebuilt Retrieval Corpora and Indexes
Rankify provides two large-scale retrieval corpora: Wikipedia (Karpukhin et al., 2020) and MS MARCO (Nguyen et al., 2017), enabling researchers to conduct retrieval experiments on well-established benchmarks. To support efficient retrieval, Rankify includes pre-builtt index files for each retriever, allowing users to directly query the corpora without requiring costly indexing. Each corpus has been indexed for multiple retrieval models, including BM25, DPR, ANCE, BGE, Contriever, and ColBERT. By providing precomputed indexes, Rankify simplifies large-scale retrieval research while maintaining consistency across different retrieval pipelines. Users can view all available datasets and retrievers using the following command (see Listing 1):
Listing 1: This script retrieves the latest information about the available datasets, and displays metadata for each dataset in the terminal.
1from rankify.dataset.dataset import Dataset
2Dataset.avaiable_dataset()
3.2. Datasets
With Rankify, we aim to provide a standardized and efficient way to handle datasets for retrieval, re-ranking, and retrieval-augmented generation (RAG). Given the complexity of retrieval workflows, Rankify integrates pre-retrieved datasets with structured annotations, allowing users to experiment with different retrieval and re-ranking methods using just a few lines of code. Our dataset design unifies multiple annotation styles and storage formats, enabling a seamless interface for loading, processing, and benchmarking retrieval-based tasks.
The datasets included in Rankify fall into multiple categories, as illustrated in Figure 3, which showcases their distribution based on task type and dataset size. Each dataset in Rankify follows a standardized schema that consists of three main components: (1) Question: The input query for which relevant information needs to be retrieved. (2) Answers: The expected correct responses associated with the question. (3) Contexts (Or Retrieved Documents): A ranked list of candidate documents retrieved by a specific retriever. The Dataset class in Rankify encapsulates these attributes and provides easy access to retrieval, evaluation, and benchmarking. The dataset schema is depicted in Figure 4, illustrating how questions, answers, and contexts interact within Rankify.
To facilitate benchmarking and ease of use, Rankify provides pre-retrieved datasets where each query is associated with 1,000 retrieved documents from different retrievers. These datasets allow users to test different re-ranking and retrieval models without the need for time-consuming indexing.
As mentioned above, to ensure standardized retrieval benchmarking, Rankify provides pre-retrieved datasets, where each query is associated with 1,000 retrieved documents from Wikipedia. These datasets span multiple retrieval tasks, including question answering, multi-hop reasoning, temporal retrieval, entity linking, fact verification, dialogue generation, and summarization. To enable fair comparisons across different retrieval methods, each dataset has been processed using multiple retrievers, including BM25, DPR, ANCE, BGE, Contriever, and ColBERT. By applying the same retrieval pipeline across all the datasets, Rankify eliminates the need for computationally expensive indexing while maintaining consistency in evaluation.
Rankify allows users to load and explore datasets with minimal effort. Below is an example demonstrating how to download and inspect a dataset using Rankify ’s API:
Listing 2: Loading a dataset in Rankify and inspecting its structure.
1from rankify.dataset.dataset import Dataset
2# Load the nq-test dataset retrieved with BM25
3dataset = Dataset(retriever="bm25", dataset_name="nq-test")
4documents = dataset.download(force_download=False)
5# Inspect the first document
6print(documents[0])
Beyond preloaded datasets, Rankify enables users to define their own datasets using the Dataset class. Below is an example demonstrating how to create a custom retrieval dataset:
Listing 3: Creating a custom dataset in Rankify.
1from rankify.dataset.dataset import Dataset, Document, Question, Answer, Context
2# Define a new dataset with a single question and answer
3question = Question("What is the capital of France?")
4answer = Answer(["Paris"])
5context = [Context(score=0.9, has_answer=True, id=1, title="France", text="The capital is Paris.")]
6# Create the document and dataset
7document = Document(question, answer, context)
8custom_dataset = Dataset(retriever="custom", dataset_name="MyDataset")
9custom_dataset.documents = [document]
10# Save the dataset
11custom_dataset.save_dataset("my_dataset.json")
Users can load their own datasets using load_dataset and load_dataset_qa functions. The load_dataset function is designed for structured datasets that include both queries and retrieved documents. On the other hand, load_dataset_qa is intended for question-answering datasets that contain only queries and answers, allowing users to use them directly for retrieval tasks. The following examples demonstrate how users can load these datasets in Rankify:
Listing 4: Loading a dataset with documents and a QA-only dataset.
1from rankify.dataset.dataset import Dataset
2# Load a dataset that includes retrieved documents
3retrieval_dataset = Dataset.load_dataset("path/to/retrieval_dataset.json")
4# Load a QA dataset with only questions and answers
5qa_dataset = Dataset.load_dataset_qa("path/to/qa_dataset.jsonl")
3.3. Retriever Models
Rankify supports a diverse set of retriever models to facilitate document retrieval for various information retrieval tasks. The current version includes both sparse and dense retrievers. Sparse retrievers, such as BM25 (Robertson et al., 1995), rely on exact term matching, while dense retrievers, such as DPR (Yates et al., 2021), ANCE (Xiong et al., 2021a), and BGE (Chen et al., 2023), and ColBERT (Khattab and Zaharia, 2020).
The retrievers implemented in Rankify are accessible via a unified interface, where users can specify the retrieval method and parameters. The available retriever models include BM25, DPR, ANCE, Contriever (Izacard et al., 2021), BGB, and ColBERT. Rankify allows users to retrieve documents efficiently using different retrieval models. The retrievers can be applied to any dataset supported by Rankify, making it flexible for experimentation and benchmarking. Users can easily initialize a retriever and apply it to a list of query-document pairs, as demonstrated in Listing 5.
Listing 5: Using a retriever model in Rankify to retrieve documents.
1from rankify.dataset.dataset import Document, Question, Answer
2from rankify.retrievers.retriever import Retriever
3
4# Define sample documents
5documents = [
6 Document(question=Question("the cast of a good day to die hard?"),
7 answers=Answer(["Jai Courtney", "Bruce Willis"]), contexts=[]),
8 Document(question=Question("Who wrote Hamlet?"),
9 answers=Answer(["Shakespeare"]), contexts=[])
10]
11# Initialize a retriever (example: ColBERT)
12retriever = Retriever(method="colbert", model="colbert-ir/colbertv2.0", n_docs=5, index_type="msmarco")
13# Retrieve documents
14retrieved_documents = retriever.retrieve(documents)
15# Print retrieved documents
16for i, doc in enumerate(retrieved_documents):
[…]
18 print(doc)
In Listing 5, users can specify the retrieval corpus, select the desired retriever model, and define the number of documents to retrieve. This flexibility allows researchers to experiment with different retrieval settings, optimizing their retrieval strategy based on the task requirements.
3.4. Re-Ranking Models
In information retrieval, two-stage pipelines are widely used to maximize retrieval performance. The first stage involves retrieving a set of candidate documents using a computationally efficient retriever, followed by a second stage where these documents are re-ranked using a stronger, typically neural network-based, model. This re-ranking step enhances retrieval quality by considering deeper semantic relationships between the query and documents. While effective, re-ranking models vary significantly in their architectures, trade-offs, and implementations, making it challenging for users to select the most suitable method for their specific needs.
To address this, Rankify provides a unified interface for re-ranking, enabling users to effortlessly switch between different models with minimal modifications. The framework supports a diverse set of 24 primary re-ranking models with 41 sub-methods, spanning different re-ranking strategies such as pointwise, pairwise, and listwise approaches. The implemented models include MonoBERT (Nogueira and Cho, 2019), MonoT5 (Nogueira et al., 2020), RankT5 (Zhuang et al., 2023), ListT5 (Yoon et al., 2024), ColBERT (Santhanam et al., 2021), RankGPT (Sun et al., 2023b), and various transformer-based re-rankers.
Users can apply these models to re-rank retrieved documents using a simple interface, as demonstrated in Listing 6.
Listing 6: Applying re-ranking in Rankify.
1from rankify.dataset.dataset import Document, Question, Answer
2from rankify.retrievers.retriever import Retriever
3from rankify.rerankers.reranker import Reranker
4# Define sample documents
5documents = [
6 Document(question=Question("the cast of a good day to die hard?"),
7 answers=Answer(["Jai Courtney", "Bruce Willis"]), contexts=[]),
8 Document(question=Question("Who wrote Hamlet?"),
9 answers=Answer(["Shakespeare"]), contexts=[])
10]
11# Initialize a retriever
12retriever = Retriever(method="colbert", model="colbert-ir/colbertv2.0", n_docs=5, index_type="msmarco")
13retrieved_documents = retriever.retrieve(documents)
14# Initialize a re-ranker
15reranker = Reranker(method="monot5",model_name="monot5-base")
16# Apply re-ranking
17reranked_documents = reranker.rerank(retrieved_documents)
18# Print re-ranked documents
19for i, doc in enumerate(reranked_documents):
[…]
21 print(doc)
Rankify supports both pre-trained models from Hugging Face and API-based re-ranking models such as RankGPT, Cohere, and Jina Reranker. Users can select from a variety of models, including cross-encoders, T5-based ranking models, and late-interaction retrieval models like ColBERT. The flexibility to switch between different re-ranking strategies allows researchers and practitioners to optimize ranking performance based on their task requirements.
3.5. Retrieval-Augmented Generation (RAG) Models
Retrieval-Augmented Generation (RAG) enhances language models by integrating retrieval mechanisms, allowing them to generate responses based on dynamically retrieved documents rather than relying solely on pre-trained knowledge. This approach is particularly effective for knowledge-intensive tasks such as open-domain question answering, fact verification, and knowledge-based text generation. Rankify provides a modular and extensible interface for applying multiple RAG methods, including zero-shot generation, Fusion-in-Decoder (FiD) (Izacard and Grave, 2020), and in-context learning  (Ram et al., 2023).
In Rankify, the Generator module enables seamless integration of RAG techniques, allowing users to experiment with different generative approaches. Users can specify the desired RAG method and model, applying generation strategies across retrieved documents.
Users can apply these methods to generate responses based on retrieved documents. Listing 7 demonstrates how to use Rankify’s RAG module with an in-context learning approach.
Listing 7: Applying Retrieval-Augmented Generation (RAG) in Rankify.
1from rankify.dataset.dataset import Document, Question, Answer, Context
2from rankify.generator.generator import Generator
3# Sample question and contexts
4question = Question("What is the capital of France?")
5answers = Answer(["Paris"])
6contexts = [
7 Context(id=1, title="France", text="The capital of France is Paris.", score=0.9),
8 Context(id=2, title="Germany", text="Berlin is the capital of Germany.", score=0.5)
9]
10# Create a Document
11doc = Document(question=question, answers=answers, contexts=contexts)
12# Initialize Generator with In-Context RALM
13generator = Generator(method="in-context-ralm", model_name=’meta-llama/Llama-3.1-8B’)
14# Generate answer
15generated_answers = generator.generate([doc, doc])
16print(generated_answers)
Rankify allows users to leverage large-scale language models such as LLaMA (Touvron et al., 2023), GPT-4 (Achiam et al., 2023), and T5-based models (Raffel et al., 2020) for retrieval-augmented generation. By supporting both encoder-decoder architectures (FiD (Izacard and Grave, 2020)) and decoder-only models (e.g., GPT, LLaMA), Rankify provides flexibility for optimizing generation quality based on task-specific requirements.
3.6. Evaluation Metrics
Rankify provides evaluation metrics for retrieval, re-ranking, and retrieval-augmented generation (RAG). For retrieval and re-ranking, Top-k accuracy measures whether relevant documents appear within the top-k results, allowing performance comparison before and after re-ranking. Listing 8 demonstrates how to compute Top-k accuracy.
Listing 8: Computing Top-k accuracy for retrieval and re-ranking in Rankify.
1from rankify.metrics.metrics import Metrics
2metrics = Metrics(documents)
3# Compute Top-k accuracy before and after re-ranking
4before_ranking = metrics.calculate_retrieval_metrics(ks=[1,5,10,20,50,100], use_reordered=False)
5after_ranking = metrics.calculate_retrieval_metrics(ks=[1,5,10,20,50,100], use_reordered=True)
6print("Before Ranking:", before_ranking)
7print("After Ranking:", after_ranking)
For QA and RAG, Rankify supports Exact Match (EM), recall, precision, and containment to evaluate generated answers. Listing 9 shows how these metrics are computed.
Listing 9: Computing QA and RAG evaluation metrics in Rankify.
1from rankify.metrics.metrics import Metrics
2qa_metrics = Metrics(documents)
3qa_results = qa_metrics.calculate_generation_metrics(generated_answers)
4print(qa_results)
By providing a unified evaluation module, Rankify ensures consistent benchmarking across retrieval, re-ranking, and RAG tasks.
4. Experimental Result and Discussion
4.1. Experiment Setup
Rankify enables researchers to benchmark retrieval, re-ranking, and RAG methods, evaluate their own approaches, and explore optimizations within these tasks. To demonstrate its capabilities, we conducted multiple experiments to provide reproducible benchmarks and performance insights. All main experiments were conducted on 8 NVIDIA A100 GPUs.
Experiments were performed on a diverse set of datasets, covering open-domain and multi-hop QA, fact verification, and temporal retrieval tasks. We evaluated retrieval performance on datasets such as Natural Questions (NQ), TriviaQA, HotpotQA, 2WikiMultiHopQA, and ArchivalQA, while re-ranking performance was analyzed on MSMARCO, WebQuestions, and PopQA. For retrieval evaluation, we measured Top-k accuracy at k=1, 5, 10, 20, 50, 100. Exact match (EM), Precision, Recall, Contains, and F1 were used as the primary evaluation metric for QA tasks.
We conducted experiments across all supported retrieval, re-ranking, and RAG methods. Retrieval models BM25, DPR, ANCE, BGE, and ColBERT. Re-ranking methods tested included MonoT5, RankT5, RankGPT, and various transformer-based re-rankers. For RAG, we evaluated zero-shot generation, Fusion-in-Decoder (FiD), in-context learning (RALM). These experiments highlight the adaptability of Rankify for benchmarking retrieval, ranking, and knowledge-grounded text generation.
Dataset Split Top-1 (%) Top-5 (%) Top-10 (%) Top-20 (%) Top-50 (%) Top-100 (%) Dataset Split Top-1 (%) Top-5 (%) Top-10 (%) Top-20 (%) Top-50 (%) Top-100 (%)
2WikiMultiHopQA (Ho et al., 2020) dev 17.17 31.72 40.14 47.73 55.88 61.17 ArchivalQA (Wang et al., 2022a) dev 18.18 33.29 39.72 45.96 53.68 58.74
train 8.85 19.78 28.15 38.16 51.16 57.70 test 17.68 32.55 39.26 45.36 52.76 58.08
ChroniclingAmericaQA (Piryani et al., 2024) test 4.18 11.23 15.83 20.87 27.49 32.48 AmbigQA (Min et al., 2020; Kwiatkowski et al., 2019) train 27.30 50.67 59.63 67.63 75.68 79.87
dev 4.26 11.80 16.42 21.50 28.19 33.15 dev 30.22 54.25 65.58 72.98 80.47 84.92
ARC (Clark et al., 2018) test 4.96 19.84 33.57 51.94 79.09 91.80 TriviaQA (Joshi et al., 2017) dev 48.70 67.57 72.84 77.38 81.19 83.59
dev 4.83 19.45 30.84 52.13 79.06 91.83 test 48.22 67.44 72.79 77.28 81.42 83.87
train 4.45 18.16 31.39 51.39 78.19 91.36 train 47.87 67.22 72.90 77.15 81.30 83.64
BoolQ (Clark et al., 2019) train 1.71 4.33 7.23 12.03 24.56 37.76 CommenseQA (Talmor et al., 2019) train 2.85 8.74 13.36 19.28 28.87 36.54
dev 1.22 3.94 6.61 11.68 23.67 37.49 dev 2.13 6.96 11.88 17.77 27.11 33.82
CuratedTREC (Lee et al., 2020) train 16.98 29.77 33.95 36.74 40.00 40.93 AY2 (Hoffart et al., 2011) train 23.81 41.60 48.84 55.88 63.78 69.62
test 18.14 31.63 34.65 38.84 41.86 43.02 dev 25.02 41.14 48.10 55.39 64.19 70.55
FEVER (Thorne et al., 2018; Petroni et al., 2021) dev 0.05 0.41 0.90 1.70 3.78 6.50 HellaSwag (Zellers et al., 2019) train 3.97 16.60 28.86 46.02 71.68 85.20
train 0.15 0.74 1.42 2.52 5.85 10.49 dev 3.77 15.00 26.31 42.89 69.00 83.51
SQuAD (Rajpurkar et al., 2016) dev 36.94 57.61 64.78 71.31 77.89 81.96 MMLU (Hendrycks et al., 2021b, a) train 3.65 14.14 23.98 38.68 62.22 78.04
test 36.67 57.37 64.50 71.09 77.74 81.84 dev 6.53 21.75 33.64 50.62 75.77 88.18
train 31.84 50.24 57.15 63.33 70.32 74.61 test 6.53 22.01 34.67 51.22 74.35 86.40
QuaRTz  (Tafjord et al., 2019) test 5.87 23.60 38.90 58.67 80.74 90.43 NarrativeQA (s Koˇciský et al., 2018) dev 15.46 23.92 27.10 30.45 35.08 38.89
dev 7.81 26.82 40.62 63.54 84.38 92.19 test 16.80 25.24 28.46 31.50 35.74 38.65
train 7.81 26.82 40.62 63.54 84.38 92.19 train 16.80 25.24 28.46 31.50 35.74 38.65
NQ (Kwiatkowski et al., 2019) dev 23.62 45.60 55.41 64.19 72.51 77.72 OpenBookQA (Mihaylov et al., 2018) dev 3.40 15.20 31.60 50.60 79.00 90.00
test 23.46 45.62 56.32 64.88 74.57 79.72 train 3.35 16.24 28.99 47.85 75.01 88.56
train 22.89 44.97 54.32 62.73 71.75 77.03 test 4.00 16.80 32.20 51.20 74.00 88.80
WebQuestions (Berant et al., 2013) test 19.54 42.86 53.44 63.24 72.34 76.43 Musique (Trivedi et al., 2022) dev 7.36 15.47 21.39 25.98 34.59 41.29
train 21.41 44.71 55.35 64.21 72.71 78.45 train 6.23 13.70 18.54 24.39 33.77 40.81
T-REx (ElSahar et al., 2018; Petroni et al., 2021) dev 45.48 64.34 70.60 76.04 80.42 82.98 TruthfulQA (Lin et al., 2022b) dev 2.33 7.59 12.12 20.69 38.19 56.55
SIQA (Sap et al., 2019) train 0.34 1.52 2.78 4.85 8.92 13.43 HotpotQA (Yang et al., 2018) dev 28.13 45.02 51.92 57.79 65.44 70.67
dev 0.61 2.05 3.58 6.55 11.31 16.58 train 34.89 51.71 58.13 64.09 71.11 75.49
StrategyQA (Geva et al., 2021) train 4.19 15.72 26.29 39.87 52.88 57.86 PopQA (Mallen et al., 2022) test 25.00 38.77 44.70 49.74 56.82 62.36
WOW (Dinan et al., 2019; Petroni et al., 2021) dev 0.20 0.33 0.36 0.39 0.46 0.46 ZSRC (Levy et al., 2017) dev 52.26 70.86 76.29 81.44 85.39 87.73
train 0.26 0.38 0.41 0.45 0.49 0.52 train 50.53 67.23 72.06 76.10 80.35 82.99
Bamboogle (Press et al., 2023) test 5.60 12.80 17.60 24.80 39.20 44.00 EntityQuestions (Sciavolino et al., 2021) test 43.36 60.50 66.08 70.61 75.41 79.06
Table 2. Retrieval performance of BM25 across several selected datasets. The table reports Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy for each dataset split. Datasets vary in complexity, including open-domain QA (e.g., TriviaQA, NQ), multi-hop reasoning (e.g., 2WikiMultiHopQA, HotpotQA), fact verification (e.g., FEVER, TruthfulQA), and temporal retrieval (e.g., ArchivalQA, ChroniclingQA).
4.2. Retrieval Results
The retrieval performance of BM25 across multiple datasets is shown in Table 2. The results report Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy for each dataset split. Performance varies significantly based on dataset characteristics, with open-domain QA datasets such as TriviaQA and Natural Questions achieving high Top-k accuracy, while more complex multi-hop datasets like 2WikiMultiHopQA and HotpotQA show lower retrieval rates, reflecting the challenges in retrieving supporting evidence across multiple documents. Temporal datasets like ArchivalQA and ChroniclingQA exhibit moderate retrieval performance, likely due to the temporal nature of their content.
For datasets requiring precise fact verification, such as FEVER and TruthfulQA, BM25 struggles, achieving low Top-1 accuracy, indicating the limitations of keyword-based retrieval in distinguishing fine-grained factual nuances. In contrast, reasoning-based datasets like StrategyQA and CommonsenseQA show relatively low retrieval accuracy, highlighting the need for more advanced retrievers capable of capturing implicit reasoning dependencies.
Table 3 presents the retrieval performance of various retrievers across three benchmark datasets: NQ, WebQ, and TriviaQA. The results report Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy, providing insights into how different retrieval models rank relevant documents. Among the tested retrievers, dense retrieval methods such as DPR, MSS-DPR, and ColBERT consistently outperform retrievers like MSS and Contriever. MSS-DPR achieves the highest Top-1 accuracy, reaching 50.16% on NQ, 44.24% on WebQ, and 61.63% on TriviaQA, demonstrating the effectiveness of fine-tuned dense retrieval. Standard DPR also performs well, particularly on NQ and TriviaQA, where it reaches 48.67% and 57.47% Top-1 accuracy, respectively. ColBERT, which employs late-interaction ranking, shows competitive performance, particularly in higher recall settings (Top-50 and Top-100).
Retriever Dataset Top-1 Top-5 Top-10 Top-20 Top-50 Top-100
MSS (Sachan et al., 2021) NQ 19.27 41.24 51.27 59.97 69.55 75.56
WebQ 11.66 29.03 39.12 49.21 61.12 68.35
TriviaQA 30.76 52.64 60.52 67.17 74.57 79.11
MSS-DPR (Sachan et al., 2021) NQ 50.16 71.88 77.47 81.44 85.98 88.14
WebQ 44.24 65.00 71.65 76.91 81.84 84.54
TriviaQA 61.63 75.21 79.14 81.85 84.92 86.58
Contriever (Izacard et al., 2021) NQ 22.16 47.28 58.72 67.86 76.01 80.55
WebQ 19.98 43.45 56.39 65.69 74.85 80.11
TriviaQA 34.16 59.48 68.00 73.91 79.83 82.93
DPR (Yates et al., 2021) NQ 48.67 68.78 74.54 79.19 83.71 85.70
WebQ 44.83 65.00 70.62 74.60 78.69 81.64
TriviaQA 57.47 72.40 76.50 79.76 82.96 85.09
ColBert (Santhanam et al., 2021) NQ 42.99 68.78 76.12 82.08 86.15 88.59
WebQ 40.11 64.81 71.56 76.57 81.20 84.50
TriviaQA 57.36 75.90 79.49 82.18 84.74 86.41
Table 3. Comparison of different retrieval models on Natural Questions (NQ), WebQuestions (WebQ), and TriviaQA. The table reports Top-1 to Top-100 accuracy for each retriever.
Reranking/ Model NQ-Test WebQ
Top-1 Top-5 Top-10 Top-20 Top-50 Top-1 Top-5 Top-10 Top-20 Top-50
BM25 - 22.10 43.76 54.45 62.93 72.74 18.80 41.83 52.16 62.40 71.70
UPR (Sachan et al., 2022) T5-small 23.60 49.44 59.97 67.70 75.15 18.55 43.65 55.56 64.91 72.68
T5-base 26.81 52.93 63.32 69.77 76.12 20.66 48.52 58.56 66.48 72.68
T5-large 29.75 55.84 65.67 72.02 76.48 24.21 51.77 60.38 67.27 73.32
T0-3B 35.42 60.44 67.56 72.68 76.75 32.48 56.39 64.17 69.09 73.67
gpt-neo-2.7B 28.75 55.01 64.81 71.13 76.56 24.75 48.91 59.64 66.83 72.63
gpt2 25.95 50.08 60.47 69.05 75.87 20.47 45.32 56.49 65.69 72.78
gpt2-medium 26.75 52.99 63.04 69.80 75.95 22.39 49.01 59.54 66.92 72.44
gpt2-large 26.59 52.35 62.68 70.13 75.95 24.06 49.85 59.84 67.07 72.78
gpt2-xl 27.28 53.85 63.24 70.11 75.84 23.57 50.78 60.48 66.87 72.73
RankGPT (Sun et al., 2023a) llamav3.1-8b 41.55 60.30 66.17 70.60 75.42 38.77 58.36 62.69 67.71 73.12
FlashRank (Damodaran, 2024) TinyBERT-L-2-v2 31.49 53.90 61.57 68.33 74.95 28.54 51.87 60.62 67.42 73.17
MiniLM-L-12-v2 41.02 62.07 67.50 72.49 76.26 37.05 58.07 64.96 70.02 73.96
MultiBERT-L-12 11.99 31.74 43.54 55.01 69.63 12.54 34.30 45.91 55.90 67.91
ce-esci-MiniLM-L12-v2 34.70 57.47 64.81 70.60 76.17 31.84 55.01 62.54 68.60 73.47
RankT5 (Zhuang et al., 2023) base 43.04 62.99 68.47 72.99 76.28 36.95 58.75 64.27 70.07 74.45
large 45.54 64.98 70.02 73.85 76.81 38.77 60.28 66.48 70.96 74.31
3b 47.17 66.45 70.85 74.18 76.89 40.40 60.33 66.58 71.16 74.45
Inranker (Laitz et al., 2024) small 15.90 40.11 46.84 56.23 69.83 14.46 36.36 46.25 57.48 69.98
base 15.90 40.69 48.11 56.62 69.66 14.46 37.45 46.80 56.54 69.68
3b 15.90 40.22 48.06 57.97 69.00 14.46 37.30 46.11 57.48 69.34
LLM2Vec (BehnamGhader et al., 2024) Meta-Llama-31-8B 24.32 49.22 59.55 68.08 75.26 26.72 50.39 60.48 68.40 73.47
MonoBert (Nogueira et al., 2019) large 39.05 61.71 67.89 72.63 76.56 34.99 57.33 64.56 70.12 73.96
Twolar (Baldelli et al., 2024) twolar-xl 46.84 65.29 70.22 73.62 76.86 41.68 60.82 67.07 71.55 74.40
Echorank (Rashid et al., 2024) flan-t5-large 36.73 55.09 59.11 62.38 62.38 31.74 53.69 58.75 61.51 61.51
flan-t5-xl 41.68 55.15 59.05 62.38 62.38 36.22 52.55 57.18 61.51 61.51
Incontext Reranker (Chen et al., 2024) llamav3.1-8b 15.15 38.50 57.11 69.66 76.48 18.89 41.83 52.16 62.40 71.70
Lit5 (Tamber et al., 2023) LiT5-Distill-base 40.05 59.80 65.95 70.80 75.73 36.76 56.79 63.48 67.81 73.12
LiT5-Distill-large 44.40 62.57 67.59 71.74 76.01 39.66 58.02 64.56 68.94 73.67
LiT5-Distill-xl 47.81 64.07 68.55 72.07 76.26 42.37 59.44 65.55 69.38 73.62
LiT5-Distill-base-v2 42.57 61.49 66.73 70.83 75.56 39.61 57.97 64.22 68.55 73.32
LiT5-Distill-large-v2 46.53 63.10 67.83 71.91 75.87 41.97 59.49 65.64 68.65 72.98
LiT5-Distill-xl-v2 47.92 64.18 69.03 72.40 76.17 41.53 60.13 65.69 69.04 73.27
Sentence Transformer Reranker all-MiniLM-L6-v2 (Wang et al., 2020) 33.35 57.34 65.37 70.72 76.01 30.95 53.44 62.10 68.01 73.52
GTR-base (Ni et al., 2021) 39.41 59.08 65.95 71.32 76.03 36.56 58.21 64.32 69.14 73.62
GTR-large 40.63 62.07 68.25 72.32 76.73 38.97 58.51 65.30 70.02 73.57
GTR-xl 41.55 61.82 67.78 72.65 76.81 38.92 59.69 66.04 70.57 74.01
GTR-xxl 42.93 63.32 68.55 73.01 77.00 39.41 60.03 65.89 70.47 74.01
T5-base (Raffel et al., 2020) 31.19 55.84 63.60 70.22 76.06 29.77 54.92 62.84 68.70 73.52
T5-xxl 38.89 61.46 67.78 72.49 76.64 35.82 58.26 65.206 70.32 74.01
T5-large 30.80 55.59 63.35 70.02 76.37 30.51 55.06 61.71 68.35 73.37
Table 4. Performance of re-ranking methods on BM25-retrieved documents for Natural Questions (NQ-Test) and WebQuestions (WebQ-Test). Results are reported in terms of Top-1, Top-5, Top-10, Top-20, and Top-50 accuracy, highlighting the impact of various re-ranking models on retrieval effectiveness.
4.3. Re-Ranking Results
The performance of various re-ranking methods applied to BM25-retrieved documents is presented in Table 4. The table reports Top-1, Top-5, Top-10, Top-20, and Top-50 accuracy for two datasets: Natural Questions (NQ-Test) and WebQuestions (WebQ). The baseline BM25 retrieval scores serve as a reference, demonstrating the extent to which re-ranking improves document ranking quality across different ranking models.
Across all models, re-ranking consistently improves retrieval effectiveness, with notable gains observed in Top-1 accuracy. Transformer-based ranking models such as UPR (T5-based models), RankGPT, FlashRank (MiniLM and TinyBERT), and RankT5 achieve substantial improvements over the BM25 baseline. In particular, RankT5-3B outperforms other ranking models, achieving a Top-1 accuracy of 47.17% on NQ-Test and 40.40% on WebQ, highlighting the effectiveness of T5-based ranking models for large-scale retrieval. RankGPT, leveraging LLaMA-3.1-8B, also demonstrates strong performance, particularly in WebQ, where it improves Top-1 accuracy from 18.80% (BM25) to 38.77%.
Among lighter-weight models, FlashRank (MiniLM and TinyBERT) and MonoBERT-large provide solid improvements while maintaining efficiency. MiniLM-L-12-v2 achieves 41.02% Top-1 accuracy on NQ-Test and 37.05% on WebQ, making it a strong candidate for deployment scenarios where computational efficiency is critical. The MonoBERT-large model reaches 39.05% on NQ-Test and 34.99% on WebQ, reinforcing the effectiveness of BERT-based cross-encoders for document ranking.
Models that integrate listwise and contrastive ranking techniques, such as Twolar, and LiT5 re-rankers, show competitive performance. Twolar-xl achieves 46.84% Top-1 accuracy on NQ-Test and 41.68% on WebQ, while LiT5-Distill-xl reaches 47.81% on NQ-Test and 42.37% on WebQ, demonstrating that fine-tuned T5 models can significantly improve ranking quality. The sentence-transformer rerankers, particularly GTR-xxl, achieve 42.93% and 39.41% Top-1 accuracy on NQ-Test and WebQ, respectively, making them effective in ranking diverse retrieval outputs.
Overall, the results highlight the importance of re-ranking in improving retrieval effectiveness. While BM25 provides a strong initial retrieval, the application of neural re-rankers significantly enhances retrieval quality, particularly for Top-1 and Top-5 results. The effectiveness of different models varies based on dataset characteristics, with larger transformer-based models excelling in ranking accuracy, while lighter-weight models offer a balance between accuracy and efficiency.
5. Conclusion
We introduced Rankify, a modular toolkit that unifies retrieval, re-ranking, and retrieval-augmented generation (RAG) within a cohesive framework. By integrating diverse retrievers, state-of-the-art re-rankers, and seamless RAG pipelines, Rankify streamlines experimentation and benchmarking in information retrieval. Our experiments demonstrate its effectiveness in enhancing ranking quality and retrieval-based text generation.
Designed for extensibility, Rankify allows for easy integration of new models and evaluation methods. Future work will focus on improving retrieval efficiency, optimizing re-ranking strategies, and advancing RAG capabilities. The framework is open-source, actively maintained, and freely available for the research community.

### v3

@@
 Abstract.
-Retrieval, re-ranking, and retrieval-augmented generation (RAG) are critical components of modern natural language processing (NLP) applications in information retrieval, question answering, and knowledge-based text generation.
+Retrieval, re-ranking, and retrieval-augmented generation (RAG) are critical components of modern applications in information retrieval, question answering, or knowledge-based text generation.
 However, existing solutions are often fragmented, lacking a unified framework that easily integrates these essential processes.
@@
 Additionally, Rankify includes a collection of pre-retrieved datasets to facilitate benchmarking, available at Huggingface [^1] https://huggingface.co/datasets/abdoelsayed/reranking-datasets-light.
-To encourage adoption and ease of integration, we provide comprehensive documentation [^2] http://rankify.readthedocs.io/, an open-source implementation on GitHub [^3] https://github.com/DataScienceUIBK/rankify, and a PyPI package for effortless installation [^4] https://pypi.org/project/rankify/.
-By providing a unified and lightweight framework, Rankify allows researchers and practitioners to advance retrieval and re-ranking methodologies while ensuring consistency, scalability, and ease of use.
+To encourage adoption and ease of integration, we provide comprehensive documentation [^2] http://rankify.readthedocs.io/, an open-source implementation on GitHub [^3] https://github.com/DataScienceUIBK/rankify, and a PyPI package for easy installation [^4] https://pypi.org/project/rankify/.
+As a unified and lightweight framework, Rankify allows researchers and practitioners to advance retrieval and re-ranking methodologies while ensuring consistency, scalability, and ease of use.
 Neural IR, Dense Retrieval, Sparse Retrieval, Re-ranking, Toolkit
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Information systems Evaluation of retrieval resultsccs: Information systems Presentation of retrieval resultsccs: Information systems Top-k retrieval in databasesccs: Information systems Rank aggregation
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; July 13–18, 2025; Padova, ITisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Information systems Evaluation of retrieval resultsccs: Information systems Presentation of retrieval resultsccs: Information systems Top-k retrieval in databasesccs: Information systems Rank aggregation
 Figure 1. Rankify logo.
 1. Introduction
-Information retrieval (IR) systems (Singhal et al., 2001; Chowdhury, 2010) are fundamental to many natural language processing (NLP) applications (Wang et al., 2024; Liu et al., 2024), including question-answering (Khamnuansin et al., 2024; Abdallah et al., 2024b), search engines (Croft et al., 2010; Xiong et al., 2024), and knowledge-based generation (Long et al., 2024).
-These systems often rely on a two-stage pipeline: a retriever that efficiently identifies a set of candidate documents and a re-ranker that refines these results to maximize relevance to the query (Yates et al., 2021; Abdallah et al., 2024a; Nogueira et al., 2019; Ren et al., 2021).
+Information retrieval (IR) systems (Singhal et al., 2001; Chowdhury, 2010) are fundamental to many applications (Wang et al., 2024; Liu et al., 2024), including question-answering (Khamnuansin et al., 2024; Abdallah et al., 2025b), search engines (Croft et al., 2010; Xiong et al., 2024), and knowledge-based generation (Long et al., 2024).
+These systems often rely on a two-stage pipeline: a retriever that efficiently identifies a set of candidate documents and a re-ranker that refines these results to maximize relevance to the query (Yates et al., 2021; Nogueira et al., 2019; Ren et al., 2021).
 This approach has proven highly effective, with retrieval and re-ranking methods achieving state-of-the-art performance across diverse NLP benchmarks.
+Table 1.
+Comparison of Retriever, Re-ranking, and RAG toolkits.
+Modular Design indicates if the toolkit uses modular components.
+Automatic Evaluation refers to the availability of built-in evaluation tools.
+Corpus shows whether the toolkit provides utilities for corpus processing, such as cleaning and chunking.
+The Combination column represents the total number of possible configurations, calculated as the product of the number of datasets, retrievers, re-rankers, and RAG methods (e.g., for Rankify: 40 × 7 × 24 × 3 = 20,160).
+Toolkit # Datasets # Retrievers # Re-rankers # RAG Methods Modular Design Automatic Evaluation Corpus Combination
+FlashRAG (Jin et al., 2024) 32 3 2 12 2,304
+FastRAG (Abane et al., 2024) 0 1 3 7 21
+AutoRAG (Kim et al., 2024) 4 2 7 1 56
+LocalRQA (Yu et al., 2024) 0 2 0 0 2
+Rerankers (Clavié, 2024) 0 0 15 0 15
+CHERCHE (Sourty et al., 2022) 0 7 2 0 14
+Rankify 40 7 24 3 20,160
 Retrievers form the backbone of information retrieval systems by identifying a subset of documents relevant to a user’s query.
@@
 Sparse retrievers, such as BM25 (Robertson et al., 1995), rely on exact term matching by representing queries and documents as high-dimensional sparse vectors.
-These methods are highly effective when query terms closely align with document content but struggle to capture semantic relationships.
-Dense retrievers, including models like DPR (Yates et al., 2021), BPR (Yamada et al., 2021), ColBERT (Santhanam et al., 2021), BGE (Chen et al., 2023), Contriever (Izacard et al., 2021) and ANCE (Xiong et al., 2021a), overcome this limitation by encoding queries and documents into low-dimensional dense vectors using neural networks, enabling retrieval based on semantic similarity even when lexical overlap is minimal.
+These methods are highly effective when query terms closely align with document content but they struggle to capture semantic relationships.
+Dense retrievers, including models like DPR (Yates et al., 2021), BPR (Yamada et al., 2021), ColBERT (Santhanam et al., 2021), BGE (Chen et al., 2023), Contriever (Izacard et al., 2021a) and ANCE (Xiong et al., 2021), overcome this limitation by encoding queries and documents into low-dimensional dense vectors using neural networks, enabling retrieval based on semantic similarity even when lexical overlap is minimal.
 However, dense retrieval requires significant computational resources for training and inference.
 Hybrid retrieval methods (Yang et al., 2019; Chen et al., 2022) integrate the strengths of both approaches, combining the precision of sparse retrievers with the semantic understanding of dense models to achieve a more balanced and robust retrieval system.
-Toolkit # Datasets # Retrievers # Re-rankers # RAG Methods Modular Design Automatic Evaluation Corpus
-FlashRAG (Jin et al., 2024) 32 3 2 12
-FastRAG (Abane et al., 2024) 0 1 3 7
-AutoRAG (Kim et al., 2024a) 4 2 7 1
-LocalRQA (Yu et al., 2024) 0 2 0 0
-Rerankers (Clavié, 2024) 0 0 15 0
-Rankify 40 7 24 3+
-Table 1.
-Comparison with other Retriever, Re-ranking, and RAG toolkits.
-Modular Component refers to whether the toolkit is composed of modular components.
-Automatic Evaluation indicates if the toolkit provides automated evaluation capabilities to assess the performance of various datasets.
-Corpus Helper shows if the toolkit offers auxiliary tools for processing corpus, including cleaning and chunking.
 Re-ranking techniques enhance the initial retrieval results by ensuring the most relevant documents appear at the top.
@@
 However, this approach treats all document pairs equally, sometimes improving lower-ranked results at the expense of top-ranked ones.
-(3) Listwise reranking (Sun et al., 2023a; Pradeep et al., 2023b; Yoon et al., 2024) considers the entire document list, prompting LLMs with the query and a subset of candidates for reranking.
+(3) Listwise reranking (Sun et al., 2023; Pradeep et al., 2023b; Yoon et al., 2024) considers the entire document list, prompting LLMs with the query and a subset of candidates for reranking.
 Due to input length limitations, these methods tend to employ a sliding window strategy, progressively refining rankings from back to front.
-Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Mozafari et al., 2024; Abdallah and Jatowt, 2023; Gao et al., 2023) enhances language models by integrating retrieval and generation, making them more effective in knowledge-intensive tasks.
+Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023) enhances language models by integrating retrieval and generation, making them more effective in knowledge-intensive tasks.
 Rather than relying solely on pre-trained knowledge, RAG dynamically retrieves relevant documents from external sources during inference, incorporating them into the generation process.
 The growing complexity and diversity of retrieval, re-ranking, and RAG methods pose significant challenges for benchmarking, reproducibility, and integration.
-Existing toolkits, such as Pyserini (Lin et al., 2021), Rerankers (Clavié, 2024) and RankLLM (Pradeep et al., 2023a), often lack flexibility, enforce rigid implementations, and require extensive preprocessing, making them less suitable for research-driven experimentation.
+Existing toolkits, such as Pyserini (Lin et al., 2021), Rerankers (Clavié, 2024) and RankLLM (Pradeep et al., 2023a) often lack flexibility, enforce rigid implementations, and require extensive preprocessing, making them less suitable for research-driven experimentation.
 Additionally, retrieval and re-ranking datasets are scattered across different sources, complicating evaluation and comparison.
-To address these challenges, we introduce Rankify, an open-source framework that unifies retrieval, re-ranking, and RAG into a modular and extensible ecosystem with its logo shown in Figure 1.
-Rankify supports diverse retrieval techniques, integrates state-of-the-art re-ranking models, and provides curated pre-retrieved datasets to streamline experimentation.
-Designed with flexibility, it enables researchers to efficiently build, evaluate, and extend retrieval pipelines while ensuring consistency in benchmarking.
+To address these challenges, we introduce Rankify, an open-source framework that unifies retrieval, re-ranking, and RAG into a modular and extensible ecosystem (logo shown in Figure 1).
+Rankify supports diverse retrieval techniques, integrates the state-of-the-art re-ranking models, and provides curated pre-retrieved datasets to streamline experimentation.
+Designed for maximizing flexibility, it enables researchers to efficiently build, evaluate, and extend retrieval pipelines while ensuring consistency in benchmarking.
 The main contributions of this paper are as follows:
-  • Curated retrieval datasets and precomputed embeddings: Rankify provides 40 datasets with 1,000 pre-retrieved documents per query across various domains (QA, dialogue, entity linking, etc.).
+  • Curated retrieval datasets and precomputed embeddings: Rankify provides 40 datasets, each with 1,000 pre-retrieved documents per query, across various domains (QA, dialogue, entity linking, etc.).
 It also includes pre-computed Wikipedia and MS MARCO corpora for multiple retrievers, eliminating preprocessing overhead.
@@
 Additionally, Rankify is freely available on PyPI and GitHub, ensuring accessibility for researchers and practitioners.
+Figure 2.
+An overview of the Rankify pipeline, demonstrating its dual capability for document retrieval.
+Users can interact with the system either by providing a query to retrieve relevant documents in real-time or by leveraging pre-retrieved datasets already indexed by the framework.
+The process starts with Pre-Retrieved Datasets & Corpus Indexing, where documents are indexed using both dense (e.g., DPR) and sparse (e.g., BM25) retrievers across corpus like Wikipedia and MS MARCO.
+Next, in the Retrieval & Re-Ranking stage, the system retrieves candidate documents using dense, or sparse retrieval methods and re-ranks them with pointwise pairwise, or listwise models powered by large language models (LLMs).
+Finally, the Generator stage applies prompt augmentation and uses models like Fusion-in-Decoder (FiD) to generate accurate and contextually informed answers.
 2. Related Work
-2.1. Retrievers
-Retrieval models have evolved significantly, transitioning from traditional sparse retrievers to sophisticated dense and hybrid retrieval techniques.
-Sparse retrievers, such as BM25 (Robertson et al., 1995), remain widely used due to their strong lexical matching capabilities and generalization.
-However, they struggle with capturing semantic relationships, leading to the rise of dense retrieval methods (Lee et al., 2019), which leverage pre-trained neural encoders to generate embeddings for queries and documents.
-Notable advancements, such as DPR (Yates et al., 2021) and ANCE (Xiong et al., 2021a), improve retrieval effectiveness by employing contrastive training with in-batch negatives (Gillick et al., 2019) and hard-negative mining (Karpukhin et al., 2020; Xiong et al., 2021b).
-Multi-vector approaches like ColBERT (Khattab and Zaharia, 2020) enable fine-grained interaction between queries and documents, albeit at the cost of increased computational overhead.
-To address the limitations of dense retrievers, hybrid models (Gao et al., 2021; Ma et al., 2022) combine sparse and dense approaches, utilizing both term matching and semantic understanding for enhanced retrieval performance.
-Recent research has also explored retrieval enhancements through knowledge distillation, curriculum learning, and pre-training.
-Distillation techniques transfer knowledge from cross-encoders or teacher retrievers to student models, improving dense retriever effectiveness (Qu et al., 2021; Chen et al., 2021).
-Curriculum learning strategies, which progressively train retrievers from easy to hard samples (Zeng et al., 2022; Lin et al., 2022a), have been shown to enhance both supervised and zero-shot retrieval effectiveness.
-Meanwhile, pre-training methods such as contrastive learning (Izacard et al., 2021) and masked autoencoding (Gao and Callan, 2021; Liu and Shao, 2022) have significantly improved the generalization of retrieval models.
-Additionally, approaches like query augmentation (Mao et al., 2021; Ma et al., 2021) generate synthetic training data to enhance retrieval robustness.
-2.2. Re-Ranking Methods
-Re-ranking plays a crucial role in refining retrieved document lists to improve ranking quality.
-Traditional approaches include pointwise learning-to-rank (LTR) methods, where each document is scored independently (Zhu et al., 2021; Cinar and Renders, 2020), and pairwise methods, which optimize the relative ordering of document pairs (Cao et al., 2007; Jia et al., 2021).
-Listwise approaches, such as LambdaRank and ListNet, directly optimize ranking metrics like Discounted Cumulative Gain (DCG) to enhance top-ranked results (Burges, 2010; Liu et al., 2017; Cao et al., 2007).
-More recently, deep neural networks and large language models (LLMs) have revolutionized re-ranking, enabling context-aware ranking adjustments.
-Cross-encoder models, which jointly encode queries and candidate documents, have demonstrated significant improvements in ranking accuracy by capturing fine-grained interactions between them.
-Additionally, zero-shot re-ranking approaches leverage models like GPT-4 for listwise re-ranking without task-specific training (Achiam et al., 2023; Sun et al., 2023c), while knowledge distillation techniques such as RankZephyr fine-tune smaller models to approximate LLM performance with reduced computational cost (Pradeep et al., 2023b).
-Re-ranking is also increasingly integrated into retrieval-augmented generation (RAG) systems, where refining retrieved documents before response generation ensures higher relevance and coherence in generated outputs.
-These advancements highlight the transition from traditional ranking techniques to sophisticated, model-driven re-ranking approaches that enhance retrieval quality in diverse NLP applications.
-Recent innovations in re-ranking include leveraging large language models (LLMs) such as GPT-4 (Achiam et al., 2023) for zero-shot and listwise re-ranking (Sun et al., 2023c).
-Other approaches, such as RankZephyr (Pradeep et al., 2023b), utilize knowledge distillation to fine-tune smaller models based on LLM outputs, enabling efficient and effective re-ranking.
-2.3. Retrieval-Augmented Generation (RAG)
-Retrieval-Augmented Generation (RAG) combines retrieval and generation to enhance language models in knowledge-intensive tasks.
-Early RAG systems (Lewis et al., 2020) retrieve context documents relevant to a query and use these as input for generation, enabling models to answer queries beyond their training data.
-More recent approaches explore the integration of retrieval with both encoder-decoder (Izacard and Grave, 2020) and decoder-only (Raffel et al., 2020) architectures, optimizing performance for a variety of downstream tasks.
-One key area of focus in RAG is mitigating hallucinations and improving the factuality of generated responses (Manakul et al., 2023).
-Techniques such as self-consistency (Wang et al., 2022b) and noisy context filtering (Fang et al., 2024) have been proposed to enhance the reliability of RAG outputs.
-However, the quality of retrieved documents remains a critical factor in the overall effectiveness of RAG pipelines.
-Recent works emphasize the importance of combining retrieval with robust re-ranking techniques to ensure that the retrieved context is both relevant and precise (Kamalloo et al., 2023).
-2.4. Frameworks
-Several frameworks have been developed to support research and development in retrieval and re-ranking systems.
-rerankers (Clavié, 2024) is a lightweight Python library that provides a unified interface to commonly used re-ranking approaches, simplifying the integration of various methods into retrieval pipelines.
-Similarly, RankLLM [^5] https://github.com/castorini/rank_llm/tree/main offers a Python toolkit focused on reproducible information retrieval research, emphasizing listwise re-ranking with large language models.
-Additionally, FlashRAG (Jin et al., 2024) and AutoRAG (Kim et al., 2024b) introduce a modular toolkit designed for efficient retrieval-augmented generation (RAG) research, aiming to enhance flexibility and extensibility in experimental settings.
-These frameworks, along with others like LangChain (Chase, 2022), LlamaIndex (Liu, 2022), and DSPy (Khattab et al., 2023), provide researchers and practitioners with versatile tools to develop and evaluate retrieval and RAG applications.
+Research in information retrieval (IR), re-ranking, and retrieval-augmented generation (RAG) has progressed significantly over the past decade.
+Traditional retrieval models like BM25 (Robertson et al., 1995) offered robust lexical matching capabilities but struggled with capturing semantic relationships.
+This limitation led to the development of dense retrieval methods (Lee et al., 2019), which leverage pre-trained neural encoders to represent queries and documents in a shared semantic space.
+Notable approaches, such as DPR (Yates et al., 2021), ANCE (Xiong et al., 2021), and multi-vector models like ColBERT (Khattab and Zaharia, 2020), have demonstrated substantial improvements in retrieval effectiveness.
+Hybrid retrievers, combining sparse and dense signals (Gao et al., 2021; Ma et al., 2022), further enhance performance by leveraging both lexical and semantic features.
+Recent advancements, including knowledge distillation (Qu et al., 2021) and curriculum learning (Zeng et al., 2022), continue to refine retrieval performance across diverse datasets.
+Re-ranking methods have also evolved alongside retrieval techniques to improve the ordering of retrieved documents.
+Traditional pointwise (Zhu et al., 2021) and pairwise (Cao et al., 2007) approaches have given way to listwise methods like LambdaRank and ListNet (Burges, 2010; Liu et al., 2017).
+Deep neural models, including cross-encoders and transformer-based architectures, have demonstrated remarkable success in re-ranking tasks by capturing complex interactions between queries and documents.
+Zero-shot and in-context re-ranking with large language models (LLMs), such as GPT-4 (Achiam et al., 2023) and RankT5 (Zhuang et al., 2023), now enable effective ranking adjustments without task-specific training.
+Retrieval-Augmented Generation (RAG) has emerged as a powerful paradigm for enhancing generative models in knowledge-intensive tasks (Lewis et al., 2020).
+By retrieving relevant documents and integrating them into the generative process, RAG systems improve factual accuracy and reduce hallucinations.
+Techniques like self-consistency (Wang et al., 2022b) and noise filtering (Fang et al., 2024) have been proposed to further improve the reliability of RAG outputs.
+However, the effectiveness of these systems heavily depends on the quality of retrieved and re-ranked documents, highlighting the need for robust retrieval frameworks.
+In response to the growing complexity of IR, re-ranking, and RAG tasks, several frameworks have been introduced.
+Rerankers (Clavié, 2024) provides a lightweight Python interface for common re-ranking models, while RankLLM focuses on listwise re-ranking with LLMs.
+Other frameworks like FlashRAG (Jin et al., 2024) and AutoRAG (Kim et al., 2024) offer modular components for RAG experimentation, though they often lack support for diverse datasets and advanced retriever configurations.
+Tools such as LangChain (Chase, 2022), LlamaIndex (Liu, 2022), and DSPy (Khattab et al., 2023) further contribute to the ecosystem by simplifying model integration and workflow design.
 A comparison of retrieval, re-ranking, and RAG toolkits is presented in Table 1, highlighting the advantages of Rankify in dataset diversity, retriever and re-ranker support, and modularity.
@@
 \$ pip install rankify
-To ensure accessibility and reproducibility, Rankify supports both precomputed and real-time retrieval workflows.
-Users can work with pre-retrieved datasets for benchmarking or perform live document retrieval and re-ranking on large-scale corpora.
+Rankify supports both precomputed and real-time retrieval workflows, allowing users to benchmark with pre-retrieved datasets or perform live document retrieval on large-scale corpora.
 The framework includes ready-to-use indices for Wikipedia and MS MARCO, eliminating the need for time-consuming indexing.
-Additionally, it supports different retrievers, giving users the flexibility to test different retrieval paradigms without modifying their pipeline.
-Rankify consists of four core modules: Datasets, Retrievers, Re-Rankers, and RAG.
-The Datasets module provides standardized access to benchmark datasets, including Natural Questions, TriviaQA, and HotpotQA, as well as the ability to create custom datasets.
-The Retrievers module integrates state-of-the-art retrieval methods, including BM25, DPR, ANCE, BGE, and ColBERT, allowing users to experiment with different retrieval strategies.
-The Re-Rankers module includes 24 ranking models with 41 sub-methods, supporting pointwise, pairwise, and listwise re-ranking techniques.
-Finally, the RAG module enables retrieval-augmented generation using models such as LLAMA, GPT, and T5, supporting in-context learning, Fusion-in-Decoder (FiD), and retrieval-augmented latent memory (RALM).
-A key advantage of Rankify is its unified API, which abstracts away implementation details, allowing users to experiment with different models and configurations using a simple interface.
-Figure 2 illustrates the architecture of Rankify, demonstrating how retrieval, ranking, and generation components interact within the pipeline.
-The framework is designed to be scalable and adaptable, making it suitable for both academic research and real-world applications in search, question answering, and document ranking.
-To encourage collaboration and community contributions, Rankify is open-source and actively maintained [^7] https://github.com/abdoelsayed2016/reranking.
-Extensive documentation [^8] http://rankify.readthedocs.io/ and inline docstrings provide clear guidance on using the framework, ensuring ease of adoption for both beginners and experienced researchers.
-The next sections explore its capabilities in greater detail, covering retrieval benchmarks, ranking evaluations, and retrieval-augmented generation results.
-Figure 3.
-Summary of datasets used in our Rankify.
-The figure illustrates the dataset sizes across different categories.
-The bubble size represents the total number of samples (Train + Dev + Test) for each dataset.
+The architecture of Rankify consists of four core modules:
+  • Datasets: Provides standardized access to datasets like Natural Questions, TriviaQA, and HotpotQA, with support for custom dataset creation.
+  • Retrievers: Integrates diverse retrieval methods such as BM25, DPR, ANCE, BGE, and ColBERT, enabling flexible retrieval strategies.
+  • Re-Rankers: Includes 24 models with 41 sub-methods, supporting pointwise, pairwise, and listwise re-ranking.
+  • RAG: Facilitates retrieval-augmented generation with models like LLAMA, GPT, and T5, supporting in-context learning, Fusion-in-Decoder (FiD), and Incontext RALM.
+A key advantage of Rankify is its unified API, which abstracts implementation details and simplifies experimentation across different models.
+Figure 2 and  3 illustrates the interactions between these modules within the pipeline.
+The framework is scalable, adaptable, and suited for both research and real-world applications like search engines and question-answering systems.
+Rankify is open-source, actively maintained [^6] https://github.com/abdoelsayed2016/reranking, and accompanied by comprehensive documentation [^7] http://rankify.readthedocs.io/ to support easy adoption for both beginners and experts.
+Table 2.
+Summary of datasets used in Rankify.
 Categories include QA, Multi-Hop QA, Long-Form QA, Multiple-Choice, Entity-Linking, Slot Filling, Fact Verification, Dialog Generation, Summarization, and Other specialized datasets.
-The log scale is used for better visualization of dataset size variations.
-Figure 4. Schema of the Dataset class, illustrating the objects used to represent a dataset in Rankify.
+Task Dataset Name # Train # Val # Test
+QA NQ (Kwiatkowski et al., 2019) 79,168 8,757 3,610
+TriviaQA (Joshi et al., 2017) 78,785 8,837 11,313
+WebQ (Berant et al., 2013) 3,778 - 2,032
+SQuAD (Rajpurkar et al., 2016) 87,599 10,570 -
+NarrativeQA (s Koˇciský et al., 2018) 32,747 3,461 10,557
+MSMARCO-QA (Nguyen et al., 2017) 808,731 101,093 -
+PopQA (Mallen et al., 2022) - - 14,267
+SIQA (Sap et al., 2019) 33,410 1,954 -
+Fermi (Kalyan et al., 2021) 8,000 1,000 1,000
+WikiQA (Yang et al., 2015) 20,360 2,733 6,165
+AmbigQA (Min et al., 2020; Kwiatkowski et al., 2019) 10,036 2,002 -
+CommenseQA (Talmor et al., 2019) 9,741 1,221 -
+PIQA (Bisk et al., 2019) 16,113 1,838 -
+BoolQ (Clark et al., 2019) 9,427 3,270 -
+Multi-Hop QA 2WikiMultiHopQA (Ho et al., 2020) 15,000 12,576 -
+Bamboogle (Press et al., 2023) - - 125
+Musique (Trivedi et al., 2022) 19,938 2,417 -
+HotpotQA (Yang et al., 2018) 90,447 7,405 -
+Temp QA ArchivalQA (Wang et al., 2022a) 384,426 48,304 48760
+ChroniclingQA (Piryani et al., 2024) 385,629 21,739 21,735
+Long-Form QA ELI5 (Fan et al., 2019) 272,634 1,507 -
+ASQA (Stelmakh et al., 2022) 4,353 948 -
+Multiple-Choice MMLU (Hendrycks et al., 2021b, a) 99,842 1,531 14,042
+TruthfulQA (Lin et al., 2022) - 817 -
+HellaSwag (Zellers et al., 2019) 39,905 10,042 -
+ARC (Clark et al., 2018) 3,370 869 3,548
+OpenBookQA (Mihaylov et al., 2018) 4,957 500 500
+Entity-linking WNED (Tedeschi et al., 2021; Petroni et al., 2021) - 8,995 -
+AIDA CoNLL-YAGO (Hoffart et al., 2011; Petroni et al., 2021) 18,395 4,784 -
+Slot filling Zero-shot RE (Levy et al., 2017; Petroni et al., 2021) 147,909 3,724 -
+T-REx (ElSahar et al., 2018; Petroni et al., 2021) 2,284,168 5,000 -
+Dialog Generation WOW (Dinan et al., 2019; Petroni et al., 2021) 63,734 3,054 -
+Fact Verification FEVER (Thorne et al., 2018; Petroni et al., 2021) 104,966 10,444 -
+Open-domain Summarization WikiAsp (Hayashi et al., 2020) 300,636 37,046 37,368
 3.1. Prebuilt Retrieval Corpora and Indexes
 Rankify provides two large-scale retrieval corpora: Wikipedia (Karpukhin et al., 2020) and MS MARCO (Nguyen et al., 2017), enabling researchers to conduct retrieval experiments on well-established benchmarks.
-To support efficient retrieval, Rankify includes pre-builtt index files for each retriever, allowing users to directly query the corpora without requiring costly indexing.
+To support efficient retrieval, Rankify includes pre-built index files for each retriever, allowing users to directly query the corpora without requiring costly indexing.
 Each corpus has been indexed for multiple retrieval models, including BM25, DPR, ANCE, BGE, Contriever, and ColBERT.
 By providing precomputed indexes, Rankify simplifies large-scale retrieval research while maintaining consistency across different retrieval pipelines.
-Users can view all available datasets and retrievers using the following command (see Listing 1):
-Listing 1: This script retrieves the latest information about the available datasets, and displays metadata for each dataset in the terminal.
+Users can view all available datasets using the following command (see Listing 1):
 1from rankify.dataset.dataset import Dataset
 2Dataset.avaiable_dataset()
+Listing 1: This script retrieves the latest information about the available datasets, and displays metadata for each dataset in the terminal.
 3.2. Datasets
-With Rankify, we aim to provide a standardized and efficient way to handle datasets for retrieval, re-ranking, and retrieval-augmented generation (RAG).
-Given the complexity of retrieval workflows, Rankify integrates pre-retrieved datasets with structured annotations, allowing users to experiment with different retrieval and re-ranking methods using just a few lines of code.
-Our dataset design unifies multiple annotation styles and storage formats, enabling a seamless interface for loading, processing, and benchmarking retrieval-based tasks.
-The datasets included in Rankify fall into multiple categories, as illustrated in Figure 3, which showcases their distribution based on task type and dataset size.
-Each dataset in Rankify follows a standardized schema that consists of three main components: (1) Question: The input query for which relevant information needs to be retrieved.
-(2) Answers: The expected correct responses associated with the question.
-(3) Contexts (Or Retrieved Documents): A ranked list of candidate documents retrieved by a specific retriever.
-The Dataset class in Rankify encapsulates these attributes and provides easy access to retrieval, evaluation, and benchmarking.
-The dataset schema is depicted in Figure 4, illustrating how questions, answers, and contexts interact within Rankify.
-To facilitate benchmarking and ease of use, Rankify provides pre-retrieved datasets where each query is associated with 1,000 retrieved documents from different retrievers.
-These datasets allow users to test different re-ranking and retrieval models without the need for time-consuming indexing.
-As mentioned above, to ensure standardized retrieval benchmarking, Rankify provides pre-retrieved datasets, where each query is associated with 1,000 retrieved documents from Wikipedia.
-These datasets span multiple retrieval tasks, including question answering, multi-hop reasoning, temporal retrieval, entity linking, fact verification, dialogue generation, and summarization.
-To enable fair comparisons across different retrieval methods, each dataset has been processed using multiple retrievers, including BM25, DPR, ANCE, BGE, Contriever, and ColBERT.
-By applying the same retrieval pipeline across all the datasets, Rankify eliminates the need for computationally expensive indexing while maintaining consistency in evaluation.
+Rankify provides a standardized and efficient way to handle datasets for retrieval, re-ranking, and retrieval-augmented generation (RAG).
+To simplify retrieval workflows, Rankify integrates pre-retrieved datasets with structured annotations, enabling users to experiment with various retrieval and re-ranking methods with minimal code.
+The datasets in Rankify are categorized based on task type and size, as shown in Table 2.
+Each dataset follows a standardized schema with three main components: the Question, which represents the query requiring relevant information; the Answers, which are the expected correct responses to the query; and the Contexts (or Retrieved Documents), which provide a ranked list of candidate documents retrieved by a specific retriever.
+To support benchmarking and reproducibility, Rankify provides pre-retrieved datasets, each with 1,000 top-ranked documents from Wikipedia [^8] Note: We are currently processing different retrievers based on MS MARCO to generate and store 1,000 top-ranked documents per query for each dataset., processed by various retrievers such as BM25, DPR, ANCE, BGE, Contriever, and ColBERT.
+These datasets cover diverse Re-Ranking and RAG tasks.
 Rankify allows users to load and explore datasets with minimal effort. Below is an example demonstrating how to download and inspect a dataset using Rankify ’s API:
-Listing 2: Loading a dataset in Rankify and inspecting its structure.
 1from rankify.dataset.dataset import Dataset
@@
 4documents = dataset.download(force_download=False)
-5# Inspect the first document
-6print(documents[0])
+Listing 2: Loading a dataset in Rankify and inspecting its structure.
 Beyond preloaded datasets, Rankify enables users to define their own datasets using the Dataset class. Below is an example demonstrating how to create a custom retrieval dataset:
+1from rankify.dataset.dataset import Dataset, Document, Question, Answer, Context
+2question = Question("What is the capital of France?")
+3answer = Answer(["Paris"])
+4context = [Context(score=0.9, has_answer=True, id=1, title="France", text="The capital is Paris.")]
+5document = Document(question, answer, context)
+6custom_dataset = Dataset(retriever="custom", dataset_name="MyDataset")
+7custom_dataset.documents = [document]
+8custom_dataset.save_dataset("my_dataset.json")
 Listing 3: Creating a custom dataset in Rankify.
-1from rankify.dataset.dataset import Dataset, Document, Question, Answer, Context
-2# Define a new dataset with a single question and answer
-3question = Question("What is the capital of France?")
-4answer = Answer(["Paris"])
-5context = [Context(score=0.9, has_answer=True, id=1, title="France", text="The capital is Paris.")]
-6# Create the document and dataset
-7document = Document(question, answer, context)
-8custom_dataset = Dataset(retriever="custom", dataset_name="MyDataset")
-9custom_dataset.documents = [document]
-10# Save the dataset
-11custom_dataset.save_dataset("my_dataset.json")
 Users can load their own datasets using load_dataset and load_dataset_qa functions.
@@
 The following examples demonstrate how users can load these datasets in Rankify:
+1from rankify.dataset.dataset import Dataset
+2retrieval_dataset = Dataset.load_dataset("path/to/retrieval_dataset.json")
+3qa_dataset = Dataset.load_dataset_qa("path/to/qa_dataset.jsonl")
 Listing 4: Loading a dataset with documents and a QA-only dataset.
-1from rankify.dataset.dataset import Dataset
-2# Load a dataset that includes retrieved documents
-3retrieval_dataset = Dataset.load_dataset("path/to/retrieval_dataset.json")
-4# Load a QA dataset with only questions and answers
-5qa_dataset = Dataset.load_dataset_qa("path/to/qa_dataset.jsonl")
 3.3. Retriever Models
 Rankify supports a diverse set of retriever models to facilitate document retrieval for various information retrieval tasks.
-The current version includes both sparse and dense retrievers.
-Sparse retrievers, such as BM25 (Robertson et al., 1995), rely on exact term matching, while dense retrievers, such as DPR (Yates et al., 2021), ANCE (Xiong et al., 2021a), and BGE (Chen et al., 2023), and ColBERT (Khattab and Zaharia, 2020).
+The current version includes both sparse retrievers, such as BM25 (Robertson et al., 1995), and dense retrievers, such as DPR (Yates et al., 2021), ANCE (Xiong et al., 2021), and BGE (Chen et al., 2023), Contriever (Izacard et al., 2021a) and ColBERT (Khattab and Zaharia, 2020).
 The retrievers implemented in Rankify are accessible via a unified interface, where users can specify the retrieval method and parameters.
-The available retriever models include BM25, DPR, ANCE, Contriever (Izacard et al., 2021), BGB, and ColBERT.
 Rankify allows users to retrieve documents efficiently using different retrieval models.
@@
 Users can easily initialize a retriever and apply it to a list of query-document pairs, as demonstrated in Listing 5.
-Listing 5: Using a retriever model in Rankify to retrieve documents.
 1from rankify.dataset.dataset import Document, Question, Answer
@@
 18 print(doc)
-In Listing 5, users can specify the retrieval corpus, select the desired retriever model, and define the number of documents to retrieve.
+Listing 5: Using a retriever model in Rankify to retrieve documents.
+In Listing 5, users can specify the retrieval corpus by setting index_type to msmarco or wiki, select the desired retriever model, and define the number of documents to retrieve.
 This flexibility allows researchers to experiment with different retrieval settings, optimizing their retrieval strategy based on the task requirements.
@@
 In information retrieval, two-stage pipelines are widely used to maximize retrieval performance.
-The first stage involves retrieving a set of candidate documents using a computationally efficient retriever, followed by a second stage where these documents are re-ranked using a stronger, typically neural network-based, model.
+The first stage involves retrieving a set of candidate documents using a computationally efficient retriever, followed by a second stage where these documents are re-ranked using a stronger, typically, neural network-based model.
 This re-ranking step enhances retrieval quality by considering deeper semantic relationships between the query and documents.
@@
 The framework supports a diverse set of 24 primary re-ranking models with 41 sub-methods, spanning different re-ranking strategies such as pointwise, pairwise, and listwise approaches.
-The implemented models include MonoBERT (Nogueira and Cho, 2019), MonoT5 (Nogueira et al., 2020), RankT5 (Zhuang et al., 2023), ListT5 (Yoon et al., 2024), ColBERT (Santhanam et al., 2021), RankGPT (Sun et al., 2023b), and various transformer-based re-rankers.
+The implemented models include MonoBERT (Nogueira and Cho, 2019), MonoT5 (Nogueira et al., 2020), RankT5 (Zhuang et al., 2023), ListT5 (Yoon et al., 2024), ColBERT (Santhanam et al., 2021), RankGPT (Sun et al., 2023), and various transformer-based re-rankers.
 Users can apply these models to re-rank retrieved documents using a simple interface, as demonstrated in Listing 6.
-Listing 6: Applying re-ranking in Rankify.
 1from rankify.dataset.dataset import Document, Question, Answer
@@
 21 print(doc)
+Listing 6: Applying re-ranking in Rankify.
 Rankify supports both pre-trained models from Hugging Face and API-based re-ranking models such as RankGPT, Cohere, and Jina Reranker.
@@
 Users can apply these methods to generate responses based on retrieved documents. Listing 7 demonstrates how to use Rankify’s RAG module with an in-context learning approach.
-Listing 7: Applying Retrieval-Augmented Generation (RAG) in Rankify.
 1from rankify.dataset.dataset import Document, Question, Answer, Context
@@
 16print(generated_answers)
+Listing 7: Applying Retrieval-Augmented Generation (RAG) in Rankify.
 Rankify allows users to leverage large-scale language models such as LLaMA (Touvron et al., 2023), GPT-4 (Achiam et al., 2023), and T5-based models (Raffel et al., 2020) for retrieval-augmented generation.
-By supporting both encoder-decoder architectures (FiD (Izacard and Grave, 2020)) and decoder-only models (e.g., GPT, LLaMA), Rankify provides flexibility for optimizing generation quality based on task-specific requirements.
+By supporting both encoder-decoder architectures (FiD (Izacard and Grave, 2020)) and decoder-only models (e.g., GPT, LLaMA), the framework provides flexibility for optimizing generation quality based on task-specific requirements.
 3.6. Evaluation Metrics
+Table 3.
+Retrieval performance of BM25 across several selected datasets.
+The table reports Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy for each dataset split.
+Datasets vary in complexity, including open-domain QA (e.g., TriviaQA, NQ), multi-hop reasoning (e.g., 2WikiMultiHopQA, HotpotQA), fact verification (e.g., TruthfulQA), and temporal retrieval (e.g., ArchivalQA, ChroniclingAmericaQA).
+Dataset Split Top-1 (%) Top-5 (%) Top-10 (%) Top-20 (%) Top-50 (%) Top-100 (%) Dataset Split Top-1 (%) Top-5 (%) Top-10 (%) Top-20 (%) Top-50 (%) Top-100 (%)
+2WikiMultiHopQA (Ho et al., 2020) Train 8.85 19.78 28.15 38.16 51.16 57.70 ArchivalQA (Wang et al., 2022a) Val 18.18 33.29 39.72 45.96 53.68 58.74
+Val 17.17 31.72 40.14 47.73 55.88 61.17 Test 17.68 32.55 39.26 45.36 52.76 58.08
+ChroniclingAmericaQA (Piryani et al., 2024) Val 4.26 11.80 16.42 21.50 28.19 33.15 AmbigQA (Min et al., 2020; Kwiatkowski et al., 2019) Train 27.30 50.67 59.63 67.63 75.68 79.87
+Test 4.18 11.23 15.83 20.87 27.49 32.48 Val 30.22 54.25 65.58 72.98 80.47 84.92
+ARC (Clark et al., 2018) Train 4.45 18.16 31.39 51.39 78.19 91.36 TriviaQA (Joshi et al., 2017) Train 47.87 67.22 72.90 77.15 81.30 83.64
+Val 4.83 19.45 30.84 52.13 79.06 91.83 Val 48.70 67.57 72.84 77.38 81.19 83.59
+Test 4.96 19.84 33.57 51.94 79.09 91.80 Test 48.22 67.44 72.79 77.28 81.42 83.87
+SQuAD (Rajpurkar et al., 2016) Train 31.84 50.24 57.15 63.33 70.32 74.61 MMLU (Hendrycks et al., 2021b, a) Train 3.65 14.14 23.98 38.68 62.22 78.04
+Val 36.94 57.61 64.78 71.31 77.89 81.96 Val 6.53 21.75 33.64 50.62 75.77 88.18
+Test 36.67 57.37 64.50 71.09 77.74 81.84 Test 6.53 22.01 34.67 51.22 74.35 86.40
+QuaRTz (Tafjord et al., 2019) Train 7.81 26.82 40.62 63.54 84.38 92.19 NarrativeQA (s Koˇciský et al., 2018) Train 16.80 25.24 28.46 31.50 35.74 38.65
+Val 7.81 26.82 40.62 63.54 84.38 92.19 Val 15.46 23.92 27.10 30.45 35.08 38.89
+Test 5.87 23.60 38.90 58.67 80.74 90.43 Test 16.80 25.24 28.46 31.50 35.74 38.65
+NQ (Kwiatkowski et al., 2019) Train 22.89 44.97 54.32 62.73 71.75 77.03 OpenBookQA (Mihaylov et al., 2018) Train 3.35 16.24 28.99 47.85 75.01 88.56
+Val 23.62 45.60 55.41 64.19 72.51 77.72 Val 3.40 15.20 31.60 50.60 79.00 90.00
+Test 23.46 45.62 56.32 64.88 74.57 79.72 Test 4.00 16.80 32.20 51.20 74.00 88.80
+WebQuestions (Berant et al., 2013) Train 21.41 44.71 55.35 64.21 72.71 78.45 Musique (Trivedi et al., 2022) Train 6.23 13.70 18.54 24.39 33.77 40.81
+Test 19.54 42.86 53.44 63.25 72.34 76.43 Val 7.36 15.47 21.39 25.98 34.59 41.29
+T-REx (ElSahar et al., 2018; Petroni et al., 2021) Val 45.48 64.34 70.60 76.04 80.42 82.98 TruthfulQA (Lin et al., 2022) Val 2.33 7.59 12.12 20.69 38.19 56.55
+SIQA (Sap et al., 2019) Train 0.34 1.52 2.78 4.85 8.92 13.43 HotpotQA (Yang et al., 2018) Train 34.89 51.71 58.13 64.09 71.11 75.49
+Val 0.61 2.05 3.58 6.55 11.31 16.58 Val 28.13 45.02 51.92 57.79 65.44 70.67
+StrategyQA (Geva et al., 2021) Train 4.19 15.72 26.29 39.87 52.88 57.86 PopQA (Mallen et al., 2022) Test 25.00 38.77 44.70 49.74 56.82 62.36
+WOW (Dinan et al., 2019; Petroni et al., 2021) Train 0.26 0.38 0.41 0.45 0.49 0.52 ZSRC (Levy et al., 2017) Train 50.53 67.23 72.06 76.10 80.35 82.99
+Val 0.20 0.33 0.36 0.39 0.46 0.46 Val 52.26 70.86 76.29 81.44 85.39 87.73
+Bamboogle (Press et al., 2023) Test 5.60 12.80 17.60 24.80 39.20 44.00 EntityQuestions (Sciavolino et al., 2021) Test 43.36 60.50 66.08 70.61 75.41 79.06
 Rankify provides evaluation metrics for retrieval, re-ranking, and retrieval-augmented generation (RAG).
-For retrieval and re-ranking, Top-k accuracy measures whether relevant documents appear within the top-k results, allowing performance comparison before and after re-ranking.
+For retrieval and re-ranking, Top-k accuracy measures whether documents containing the correct answer appear within the top-k results.
+This definition of relevance, while task-specific, aligns with the goal of downstream applications like question answering.
 Listing 8 demonstrates how to compute Top-k accuracy.
-Listing 8: Computing Top-k accuracy for retrieval and re-ranking in Rankify.
 1from rankify.metrics.metrics import Metrics
@@
 7print("After Ranking:", after_ranking)
+Listing 8: Computing Top-k accuracy for retrieval and re-ranking in Rankify.
 For QA and RAG, Rankify supports Exact Match (EM), recall, precision, and containment to evaluate generated answers. Listing 9 shows how these metrics are computed.
-Listing 9: Computing QA and RAG evaluation metrics in Rankify.
 1from rankify.metrics.metrics import Metrics
@@
 4print(qa_results)
+Listing 9: Computing QA and RAG evaluation metrics in Rankify.
 By providing a unified evaluation module, Rankify ensures consistent benchmarking across retrieval, re-ranking, and RAG tasks.
@@
 To demonstrate its capabilities, we conducted multiple experiments to provide reproducible benchmarks and performance insights.
-All main experiments were conducted on 8 NVIDIA A100 GPUs.
+All main experiments were conducted on 2x NVIDIA A100 GPUs.
 Experiments were performed on a diverse set of datasets, covering open-domain and multi-hop QA, fact verification, and temporal retrieval tasks.
+All the experiments in our study use preprocessed English Wikipedia dump from December 2018 as released by (Yates et al., 2021) as evidence passages.
+Each Wikipedia article is split into non-overlapping 100-word passages, with over 21 million passages in total.
 We evaluated retrieval performance on datasets such as Natural Questions (NQ), TriviaQA, HotpotQA, 2WikiMultiHopQA, and ArchivalQA, while re-ranking performance was analyzed on MSMARCO, WebQuestions, and PopQA.
 For retrieval evaluation, we measured Top-k accuracy at k=1, 5, 10, 20, 50, 100.
-Exact match (EM), Precision, Recall, Contains, and F1 were used as the primary evaluation metric for QA tasks.
+Exact match (EM), Precision, Recall, Contains, and F1 were used as the primary evaluation metrics for QA tasks.
 We conducted experiments across all supported retrieval, re-ranking, and RAG methods.
@@
 These experiments highlight the adaptability of Rankify for benchmarking retrieval, ranking, and knowledge-grounded text generation.
-Dataset Split Top-1 (%) Top-5 (%) Top-10 (%) Top-20 (%) Top-50 (%) Top-100 (%) Dataset Split Top-1 (%) Top-5 (%) Top-10 (%) Top-20 (%) Top-50 (%) Top-100 (%)
-2WikiMultiHopQA (Ho et al., 2020) dev 17.17 31.72 40.14 47.73 55.88 61.17 ArchivalQA (Wang et al., 2022a) dev 18.18 33.29 39.72 45.96 53.68 58.74
-train 8.85 19.78 28.15 38.16 51.16 57.70 test 17.68 32.55 39.26 45.36 52.76 58.08
-ChroniclingAmericaQA (Piryani et al., 2024) test 4.18 11.23 15.83 20.87 27.49 32.48 AmbigQA (Min et al., 2020; Kwiatkowski et al., 2019) train 27.30 50.67 59.63 67.63 75.68 79.87
-dev 4.26 11.80 16.42 21.50 28.19 33.15 dev 30.22 54.25 65.58 72.98 80.47 84.92
-ARC (Clark et al., 2018) test 4.96 19.84 33.57 51.94 79.09 91.80 TriviaQA (Joshi et al., 2017) dev 48.70 67.57 72.84 77.38 81.19 83.59
-dev 4.83 19.45 30.84 52.13 79.06 91.83 test 48.22 67.44 72.79 77.28 81.42 83.87
-train 4.45 18.16 31.39 51.39 78.19 91.36 train 47.87 67.22 72.90 77.15 81.30 83.64
-BoolQ (Clark et al., 2019) train 1.71 4.33 7.23 12.03 24.56 37.76 CommenseQA (Talmor et al., 2019) train 2.85 8.74 13.36 19.28 28.87 36.54
-dev 1.22 3.94 6.61 11.68 23.67 37.49 dev 2.13 6.96 11.88 17.77 27.11 33.82
-CuratedTREC (Lee et al., 2020) train 16.98 29.77 33.95 36.74 40.00 40.93 AY2 (Hoffart et al., 2011) train 23.81 41.60 48.84 55.88 63.78 69.62
-test 18.14 31.63 34.65 38.84 41.86 43.02 dev 25.02 41.14 48.10 55.39 64.19 70.55
-FEVER (Thorne et al., 2018; Petroni et al., 2021) dev 0.05 0.41 0.90 1.70 3.78 6.50 HellaSwag (Zellers et al., 2019) train 3.97 16.60 28.86 46.02 71.68 85.20
-train 0.15 0.74 1.42 2.52 5.85 10.49 dev 3.77 15.00 26.31 42.89 69.00 83.51
-SQuAD (Rajpurkar et al., 2016) dev 36.94 57.61 64.78 71.31 77.89 81.96 MMLU (Hendrycks et al., 2021b, a) train 3.65 14.14 23.98 38.68 62.22 78.04
-test 36.67 57.37 64.50 71.09 77.74 81.84 dev 6.53 21.75 33.64 50.62 75.77 88.18
-train 31.84 50.24 57.15 63.33 70.32 74.61 test 6.53 22.01 34.67 51.22 74.35 86.40
-QuaRTz  (Tafjord et al., 2019) test 5.87 23.60 38.90 58.67 80.74 90.43 NarrativeQA (s Koˇciský et al., 2018) dev 15.46 23.92 27.10 30.45 35.08 38.89
-dev 7.81 26.82 40.62 63.54 84.38 92.19 test 16.80 25.24 28.46 31.50 35.74 38.65
-train 7.81 26.82 40.62 63.54 84.38 92.19 train 16.80 25.24 28.46 31.50 35.74 38.65
-NQ (Kwiatkowski et al., 2019) dev 23.62 45.60 55.41 64.19 72.51 77.72 OpenBookQA (Mihaylov et al., 2018) dev 3.40 15.20 31.60 50.60 79.00 90.00
-test 23.46 45.62 56.32 64.88 74.57 79.72 train 3.35 16.24 28.99 47.85 75.01 88.56
-train 22.89 44.97 54.32 62.73 71.75 77.03 test 4.00 16.80 32.20 51.20 74.00 88.80
-WebQuestions (Berant et al., 2013) test 19.54 42.86 53.44 63.24 72.34 76.43 Musique (Trivedi et al., 2022) dev 7.36 15.47 21.39 25.98 34.59 41.29
-train 21.41 44.71 55.35 64.21 72.71 78.45 train 6.23 13.70 18.54 24.39 33.77 40.81
-T-REx (ElSahar et al., 2018; Petroni et al., 2021) dev 45.48 64.34 70.60 76.04 80.42 82.98 TruthfulQA (Lin et al., 2022b) dev 2.33 7.59 12.12 20.69 38.19 56.55
-SIQA (Sap et al., 2019) train 0.34 1.52 2.78 4.85 8.92 13.43 HotpotQA (Yang et al., 2018) dev 28.13 45.02 51.92 57.79 65.44 70.67
-dev 0.61 2.05 3.58 6.55 11.31 16.58 train 34.89 51.71 58.13 64.09 71.11 75.49
-StrategyQA (Geva et al., 2021) train 4.19 15.72 26.29 39.87 52.88 57.86 PopQA (Mallen et al., 2022) test 25.00 38.77 44.70 49.74 56.82 62.36
-WOW (Dinan et al., 2019; Petroni et al., 2021) dev 0.20 0.33 0.36 0.39 0.46 0.46 ZSRC (Levy et al., 2017) dev 52.26 70.86 76.29 81.44 85.39 87.73
-train 0.26 0.38 0.41 0.45 0.49 0.52 train 50.53 67.23 72.06 76.10 80.35 82.99
-Bamboogle (Press et al., 2023) test 5.60 12.80 17.60 24.80 39.20 44.00 EntityQuestions (Sciavolino et al., 2021) test 43.36 60.50 66.08 70.61 75.41 79.06
+Table 4. Comparison of different retrieval models on NQ, WebQ, and TriviaQA. The table reports Top-1 to Top-100 accuracy for each retriever.
+Retriever Dataset Top-1 Top-5 Top-10 Top-20 Top-50 Top-100
+Contriever (Izacard et al., 2021a) NQ 38.81 65.65 73.91 79.56 84.88 88.01
+WebQ 35.97 63.83 69.78 75.94 81.64 83.76
+TriviaQA 49.85 71.39 76.68 80.26 83.85 85.72
+DPR (Yates et al., 2021) NQ 44.57 67.76 74.52 79.50 84.40 86.81
+WebQ 44.64 63.98 70.52 75.05 80.51 82.97
+TriviaQA 57.47 72.40 76.50 79.76 82.96 85.09
+ColBert (Santhanam et al., 2021) NQ 42.99 68.78 76.12 82.08 86.15 88.59
+WebQ 40.11 64.81 71.56 76.57 81.20 84.50
+TriviaQA 57.36 75.90 79.49 82.18 84.74 86.41
+Ance (Xiong et al., 2021) NQ 50.80 71.86 78.12 82.52 86.23 88.28
+WebQ 46.51 66.24 71.80 76.62 81.79 84.25
+TriviaQA 56.46 72.14 76.65 80.04 83.37 85.22
+BGE (Xiao et al., 2023; Chen et al., 2023) NQ 48.03 72.22 78.50 82.66 86.90 89.45
+WebQ 42.67 65.45 72.54 79.04 82.78 85.88
+TriviaQA 57.81 75.39 79.83 82.92 85.53 86.94
 Table 5.
-Retrieval performance of BM25 across several selected datasets.
-The table reports Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy for each dataset split.
-Datasets vary in complexity, including open-domain QA (e.g., TriviaQA, NQ), multi-hop reasoning (e.g., 2WikiMultiHopQA, HotpotQA), fact verification (e.g., FEVER, TruthfulQA), and temporal retrieval (e.g., ArchivalQA, ChroniclingQA).
+Top-20/Top-100 retrieval accuracy of different retrievers on three open-domain QA datasets (NQ, TriviaQA, and WebQ).
+Our Rankify implementation achieves results identical to Pyserini across all retrievers, as we use Pyserini as the indexing backend.
+This validates the correctness of our implementation.
+We also match the official GitHub results for ColBERT and Contriever on these datasets.
+Retrievers NQ TriviaQA WebQ
+DPR (Yates et al., 2021) 78.4 / 85.4 79.4 / 85.0 73.2 / 81.4
+DPR_Pyserini (Lin et al., 2021) 79.5 / 86.8 79.7 / 85.1 75.1 / 82.9
+DPR_Rankify 79.5 / 86.8 79.7 / 85.1 75.1 / 82.9
+BM25 (Sachan et al., 2022) 63.0 / 78.2 76.4 / 83.1 62.3 / 75.5
+BM25_Pyserini (Lin et al., 2021) 64.8 / 79.7 77.3 / 83.8 63.3 / 76.4
+BM25_Rankify 64.8 / 79.7 77.3 / 83.8 63.3 / 76.4
+Contriever (Izacard et al., 2021b) 79.6 / 88.0 80.4 / 85.7 75.9 / 83.7
+Contriever_Rankify 79.6 / 88.0 80.3 / 85.7 75.9 / 83.7
+ColBert (Santhanam et al., 2021) 82.1 / 88.5 82.2 / 86.4 76.6 / 84.5
+ColBert_Rankify 82.1 / 88.5 82.2 / 86.4 76.6 / 84.5
+ANCE (Xiong et al., 2021) 82.1 / 87.9 80.3 / 85.2 - / -
+ANCE_Rankify 82.5 / 88.5 80.0 / 85.2 76.6 / 84.25
+Table 6.
+Performance of re-ranking methods on BM25-retrieved documents for NQ Test and WebQ Test.
+Results are reported in terms of Top-1, Top-5, Top-10, Top-20, and Top-50 accuracy, highlighting the impact of various re-ranking models on retrieval effectiveness.
+Please note that some results may differ from the original papers (e.g., UPR) as our experiments were conducted with the top 100 retrieved documents, whereas the original studies used 1,000 documents for ranking.
+Reranking/ Model NQ WebQ
+Top-1 Top-10 Top-50 Top-1 Top-10 Top-50
+BM25 - 23.46 56.32 74.57 19.54 53.44 72.34
+UPR (Sachan et al., 2022) T5-small 23.60 59.97 75.15 18.55 55.56 72.68
+T5-base 26.81 63.32 76.12 20.66 58.56 72.68
+T5-large 29.75 65.67 76.48 24.21 60.38 73.32
+T0-3B 35.42 67.56 76.75 32.48 64.17 73.67
+gpt2 25.95 60.47 75.87 20.47 56.49 72.78
+gpt2-medium 26.75 63.04 75.95 22.39 59.54 72.44
+gpt2-large 26.59 62.68 75.95 24.06 59.84 72.78
+gpt2-xl 27.28 63.24 75.84 23.57 60.48 72.73
+gpt-neo-2.7B 28.75 64.81 76.56 24.75 59.64 72.63
+RankGPT (Sun et al., 2023) llamav3.1-8b 41.55 66.17 75.42 38.77 62.69 73.12
+FlashRank (Damodaran, 2024) TinyBERT-L-2-v2 31.49 61.57 74.95 28.54 60.62 73.17
+MultiBERT-L-12 11.99 43.54 69.63 12.54 45.91 67.91
+ce-esci-MiniLM-L12-v2 34.70 64.81 76.17 31.84 62.54 73.47
+T5-flan 7.95 36.14 66.67 12.05 42.96 67.27
+RankT5 (Zhuang et al., 2023) base 43.04 68.47 76.28 36.95 64.27 74.45
+large 45.54 70.02 76.81 38.77 66.48 74.31
+3b 47.17 70.85 76.89 40.40 66.58 74.45
+Inranker (Laitz et al., 2024) small 15.90 46.84 69.83 14.46 46.25 69.98
+base 15.90 48.11 69.66 14.46 46.80 69.68
+3b 15.90 48.06 69.00 14.46 46.11 69.34
+LLM2Vec (BehnamGhader et al., 2024) Meta-Llama-31-8B 24.32 59.55 75.26 26.72 60.48 73.47
+MonoBert (Nogueira et al., 2019) large 39.05 67.89 76.56 34.99 64.56 73.96
+Twolar (Baldelli et al., 2024) twolar-xl 46.84 70.22 76.86 41.68 67.07 74.40
+Echorank (Rashid et al., 2024) flan-t5-large 36.73 59.11 62.38 31.74 58.75 61.51
+flan-t5-xl 41.68 59.05 62.38 36.22 57.18 61.51
+Incontext Reranker (Chen et al., 2024) llamav3.1-8b 15.15 57.11 76.48 18.89 52.16 71.70
+Lit5 (Tamber et al., 2023) LiT5-Distill-base 40.05 65.95 75.73 36.76 63.48 73.12
+LiT5-Distill-large 44.40 67.59 76.01 39.66 64.56 73.67
+LiT5-Distill-xl 47.81 68.55 76.26 42.37 65.55 73.62
+LiT5-Distill-base-v2 42.57 66.73 75.56 39.61 64.22 73.32
+LiT5-Distill-large-v2 46.53 67.83 75.87 41.97 65.64 72.98
+LiT5-Distill-xl-v2 47.92 69.03 76.17 41.53 65.69 73.27
+Sentence Transformer Reranker GTR-base (Ni et al., 2021) 39.41 65.95 76.03 36.56 64.32 73.62
+GTR-large 40.63 68.25 76.73 38.97 65.30 73.57
+T5-base (Raffel et al., 2020) 31.19 63.60 76.06 29.77 62.84 73.52
+T5-large 30.80 63.35 76.37 30.51 61.71 73.37
+all-MiniLM-L6-v2 (Wang et al., 2020) 33.35 65.37 76.01 30.95 62.10 73.52
+GTR-xl 41.55 67.78 76.81 38.92 66.04 74.01
+GTR-xxl 42.93 68.55 77.00 39.41 65.89 74.01
+T5-xxl 38.89 67.78 76.64 35.82 65.20 74.01
+Bert-co-condensor 30.96 61.91 75.20 32.43 62.20 73.08
+Roberta-base-v2 32.60 63.24 75.42 31.34 62.64 73.37
 4.2. Retrieval Results
@@
 The results report Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy for each dataset split.
-Performance varies significantly based on dataset characteristics, with open-domain QA datasets such as TriviaQA and Natural Questions achieving high Top-k accuracy, while more complex multi-hop datasets like 2WikiMultiHopQA and HotpotQA show lower retrieval rates, reflecting the challenges in retrieving supporting evidence across multiple documents.
-Temporal datasets like ArchivalQA and ChroniclingQA exhibit moderate retrieval performance, likely due to the temporal nature of their content.
-For datasets requiring precise fact verification, such as FEVER and TruthfulQA, BM25 struggles, achieving low Top-1 accuracy, indicating the limitations of keyword-based retrieval in distinguishing fine-grained factual nuances.
-In contrast, reasoning-based datasets like StrategyQA and CommonsenseQA show relatively low retrieval accuracy, highlighting the need for more advanced retrievers capable of capturing implicit reasoning dependencies.
-Table 3 presents the retrieval performance of various retrievers across three benchmark datasets: NQ, WebQ, and TriviaQA.
+Performance varies significantly based on dataset characteristics, with high Top-k accuracy achieved on open-domain QA datasets such as TriviaQA and Natural Questions.
+On the other hand, more complex multi-hop datasets like 2WikiMultiHopQA and HotpotQA are characterized by lower retrieval rates, reflecting the challenges in retrieving supporting evidence across multiple documents.
+Temporal datasets like ArchivalQA and ChroniclingAmericaQA exhibit moderate retrieval performance, likely due to the temporal nature of their content.
+It is important to note that some datasets in Table 3 were either not officially recorded in previous studies or rely on different underlying corpora, making direct comparison difficult.
+Additionally, some datasets (e.g., ChroniclingAmericaQA) do not report retrieval accuracy as a standard Top-k metric, further complicating direct validation.
+Despite these variations, our results shown in Table 3 provide a comprehensive overview of BM25’s strengths and limitations.
+Table 4 presents the retrieval performance of various dense retrievers across three benchmark datasets: NQ, WebQ, and TriviaQA.
 The results report Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy, providing insights into how different retrieval models rank relevant documents.
@@
 ColBERT, which employs late-interaction ranking, shows competitive performance, particularly in higher recall settings (Top-50 and Top-100).
-Retriever Dataset Top-1 Top-5 Top-10 Top-20 Top-50 Top-100
-MSS (Sachan et al., 2021) NQ 19.27 41.24 51.27 59.97 69.55 75.56
-WebQ 11.66 29.03 39.12 49.21 61.12 68.35
-TriviaQA 30.76 52.64 60.52 67.17 74.57 79.11
-MSS-DPR (Sachan et al., 2021) NQ 50.16 71.88 77.47 81.44 85.98 88.14
-WebQ 44.24 65.00 71.65 76.91 81.84 84.54
-TriviaQA 61.63 75.21 79.14 81.85 84.92 86.58
-Contriever (Izacard et al., 2021) NQ 22.16 47.28 58.72 67.86 76.01 80.55
-WebQ 19.98 43.45 56.39 65.69 74.85 80.11
-TriviaQA 34.16 59.48 68.00 73.91 79.83 82.93
-DPR (Yates et al., 2021) NQ 48.67 68.78 74.54 79.19 83.71 85.70
-WebQ 44.83 65.00 70.62 74.60 78.69 81.64
-TriviaQA 57.47 72.40 76.50 79.76 82.96 85.09
-ColBert (Santhanam et al., 2021) NQ 42.99 68.78 76.12 82.08 86.15 88.59
-WebQ 40.11 64.81 71.56 76.57 81.20 84.50
-TriviaQA 57.36 75.90 79.49 82.18 84.74 86.41
-Table 3. Comparison of different retrieval models on Natural Questions (NQ), WebQuestions (WebQ), and TriviaQA. The table reports Top-1 to Top-100 accuracy for each retriever.
-Reranking/ Model NQ-Test WebQ
-Top-1 Top-5 Top-10 Top-20 Top-50 Top-1 Top-5 Top-10 Top-20 Top-50
-BM25 - 22.10 43.76 54.45 62.93 72.74 18.80 41.83 52.16 62.40 71.70
-UPR (Sachan et al., 2022) T5-small 23.60 49.44 59.97 67.70 75.15 18.55 43.65 55.56 64.91 72.68
-T5-base 26.81 52.93 63.32 69.77 76.12 20.66 48.52 58.56 66.48 72.68
-T5-large 29.75 55.84 65.67 72.02 76.48 24.21 51.77 60.38 67.27 73.32
-T0-3B 35.42 60.44 67.56 72.68 76.75 32.48 56.39 64.17 69.09 73.67
-gpt-neo-2.7B 28.75 55.01 64.81 71.13 76.56 24.75 48.91 59.64 66.83 72.63
-gpt2 25.95 50.08 60.47 69.05 75.87 20.47 45.32 56.49 65.69 72.78
-gpt2-medium 26.75 52.99 63.04 69.80 75.95 22.39 49.01 59.54 66.92 72.44
-gpt2-large 26.59 52.35 62.68 70.13 75.95 24.06 49.85 59.84 67.07 72.78
-gpt2-xl 27.28 53.85 63.24 70.11 75.84 23.57 50.78 60.48 66.87 72.73
-RankGPT (Sun et al., 2023a) llamav3.1-8b 41.55 60.30 66.17 70.60 75.42 38.77 58.36 62.69 67.71 73.12
-FlashRank (Damodaran, 2024) TinyBERT-L-2-v2 31.49 53.90 61.57 68.33 74.95 28.54 51.87 60.62 67.42 73.17
-MiniLM-L-12-v2 41.02 62.07 67.50 72.49 76.26 37.05 58.07 64.96 70.02 73.96
-MultiBERT-L-12 11.99 31.74 43.54 55.01 69.63 12.54 34.30 45.91 55.90 67.91
-ce-esci-MiniLM-L12-v2 34.70 57.47 64.81 70.60 76.17 31.84 55.01 62.54 68.60 73.47
-RankT5 (Zhuang et al., 2023) base 43.04 62.99 68.47 72.99 76.28 36.95 58.75 64.27 70.07 74.45
-large 45.54 64.98 70.02 73.85 76.81 38.77 60.28 66.48 70.96 74.31
-3b 47.17 66.45 70.85 74.18 76.89 40.40 60.33 66.58 71.16 74.45
-Inranker (Laitz et al., 2024) small 15.90 40.11 46.84 56.23 69.83 14.46 36.36 46.25 57.48 69.98
-base 15.90 40.69 48.11 56.62 69.66 14.46 37.45 46.80 56.54 69.68
-3b 15.90 40.22 48.06 57.97 69.00 14.46 37.30 46.11 57.48 69.34
-LLM2Vec (BehnamGhader et al., 2024) Meta-Llama-31-8B 24.32 49.22 59.55 68.08 75.26 26.72 50.39 60.48 68.40 73.47
-MonoBert (Nogueira et al., 2019) large 39.05 61.71 67.89 72.63 76.56 34.99 57.33 64.56 70.12 73.96
-Twolar (Baldelli et al., 2024) twolar-xl 46.84 65.29 70.22 73.62 76.86 41.68 60.82 67.07 71.55 74.40
-Echorank (Rashid et al., 2024) flan-t5-large 36.73 55.09 59.11 62.38 62.38 31.74 53.69 58.75 61.51 61.51
-flan-t5-xl 41.68 55.15 59.05 62.38 62.38 36.22 52.55 57.18 61.51 61.51
-Incontext Reranker (Chen et al., 2024) llamav3.1-8b 15.15 38.50 57.11 69.66 76.48 18.89 41.83 52.16 62.40 71.70
-Lit5 (Tamber et al., 2023) LiT5-Distill-base 40.05 59.80 65.95 70.80 75.73 36.76 56.79 63.48 67.81 73.12
-LiT5-Distill-large 44.40 62.57 67.59 71.74 76.01 39.66 58.02 64.56 68.94 73.67
-LiT5-Distill-xl 47.81 64.07 68.55 72.07 76.26 42.37 59.44 65.55 69.38 73.62
-LiT5-Distill-base-v2 42.57 61.49 66.73 70.83 75.56 39.61 57.97 64.22 68.55 73.32
-LiT5-Distill-large-v2 46.53 63.10 67.83 71.91 75.87 41.97 59.49 65.64 68.65 72.98
-LiT5-Distill-xl-v2 47.92 64.18 69.03 72.40 76.17 41.53 60.13 65.69 69.04 73.27
-Sentence Transformer Reranker all-MiniLM-L6-v2 (Wang et al., 2020) 33.35 57.34 65.37 70.72 76.01 30.95 53.44 62.10 68.01 73.52
-GTR-base (Ni et al., 2021) 39.41 59.08 65.95 71.32 76.03 36.56 58.21 64.32 69.14 73.62
-GTR-large 40.63 62.07 68.25 72.32 76.73 38.97 58.51 65.30 70.02 73.57
-GTR-xl 41.55 61.82 67.78 72.65 76.81 38.92 59.69 66.04 70.57 74.01
-GTR-xxl 42.93 63.32 68.55 73.01 77.00 39.41 60.03 65.89 70.47 74.01
-T5-base (Raffel et al., 2020) 31.19 55.84 63.60 70.22 76.06 29.77 54.92 62.84 68.70 73.52
-T5-xxl 38.89 61.46 67.78 72.49 76.64 35.82 58.26 65.206 70.32 74.01
-T5-large 30.80 55.59 63.35 70.02 76.37 30.51 55.06 61.71 68.35 73.37
-Table 4.
-Performance of re-ranking methods on BM25-retrieved documents for Natural Questions (NQ-Test) and WebQuestions (WebQ-Test).
-Results are reported in terms of Top-1, Top-5, Top-10, Top-20, and Top-50 accuracy, highlighting the impact of various re-ranking models on retrieval effectiveness.
+Figure 4. Exact Match (EM) for BM25, Contriever, and DPR retrievers across three datasets (NQ, TriviaQA, WebQ) using various language models (LLaMA V3/V3.1, Gemma 2B/9B, LLaMA 2 13B, and Mistral 7B).
+4.3. Comparison with Original Implementations
+To validate the correctness and reliability of the Rankify’s retrieval module, we compare its performance against original implementations and Pyserini-based baselines for multiple retrievers.
+Specifically, we evaluate DPR, Contriever, BM25, and ColBERT across three open-domain QA datasets: NQ, TriviaQA, and WebQ.
+The retrieval performance is measured using Top-20 and Top-100 accuracy.
+Table 5 presents the results of our Rankify implementation alongside the original models and Pyserini-based implementations.
+Our results show that Rankify achieves retrieval performance identical to Pyserini for DPR, BM25 as we utilize Pyserini as the backend for indexing and retrieval.
+However, for Contriever BGE, and ColBERT, we implemented their methods independently, and our results closely match the original implementations, demonstrating the correctness of our approach.
+Additionally, we observe improvements in the performance of BM25 and DPR compared to their original baselines.
+These improvements can be attributed to the optimizations present in Pyserini, such as better indexing techniques and parameter tuning.
+For ColBERT and Contriever, we replicate the results reported in their respective papers by running their official GitHub implementations on our datasets, yielding identical outcomes.
 4.4. Re-Ranking Results
 The performance of various re-ranking methods applied to BM25-retrieved documents is presented in Table 6.
-The table reports Top-1, Top-5, Top-10, Top-20, and Top-50 accuracy for two datasets: Natural Questions (NQ-Test) and WebQuestions (WebQ).
-The baseline BM25 retrieval scores serve as a reference, demonstrating the extent to which re-ranking improves document ranking quality across different ranking models.
-Across all models, re-ranking consistently improves retrieval effectiveness, with notable gains observed in Top-1 accuracy.
-Transformer-based ranking models such as UPR (T5-based models), RankGPT, FlashRank (MiniLM and TinyBERT), and RankT5 achieve substantial improvements over the BM25 baseline.
-In particular, RankT5-3B outperforms other ranking models, achieving a Top-1 accuracy of 47.17% on NQ-Test and 40.40% on WebQ, highlighting the effectiveness of T5-based ranking models for large-scale retrieval.
-RankGPT, leveraging LLaMA-3.1-8B, also demonstrates strong performance, particularly in WebQ, where it improves Top-1 accuracy from 18.80% (BM25) to 38.77%.
+The table reports Top-1, Top-10, and Top-50 accuracy for the NQ-Test and WebQ datasets.
+The baseline BM25 retrieval scores serve as a reference, demonstrating the extent to which re-ranking improves document ranking quality across different models.
+Across all methods, re-ranking consistently improves retrieval effectiveness, with notable gains observed in Top-1 accuracy.
+Transformer-based models such as UPR (T5-based), RankGPT, FlashRank (MiniLM and TinyBERT), and RankT5 achieve substantial improvements over the BM25 baseline.
+In particular, RankT5-3B outperforms other models, achieving a Top-1 accuracy of 47.17% on NQ-Test and 40.40% on WebQ, highlighting the effectiveness of T5-based models for large-scale retrieval.
+RankGPT, leveraging LLaMA-3.1-8B, also shows strong performance, particularly in WebQ, where it improves Top-1 accuracy from 18.80% (BM25) to 38.77%.
 Among lighter-weight models, FlashRank (MiniLM and TinyBERT) and MonoBERT-large provide solid improvements while maintaining efficiency.
@@
 The MonoBERT-large model reaches 39.05% on NQ-Test and 34.99% on WebQ, reinforcing the effectiveness of BERT-based cross-encoders for document ranking.
-Models that integrate listwise and contrastive ranking techniques, such as Twolar, and LiT5 re-rankers, show competitive performance.
+Models that integrate listwise and contrastive ranking techniques, such as Twolar-xl and LiT5, show competitive performance.
 Twolar-xl achieves 46.84% Top-1 accuracy on NQ-Test and 41.68% on WebQ, while LiT5-Distill-xl reaches 47.81% on NQ-Test and 42.37% on WebQ, demonstrating that fine-tuned T5 models can significantly improve ranking quality.
-The sentence-transformer rerankers, particularly GTR-xxl, achieve 42.93% and 39.41% Top-1 accuracy on NQ-Test and WebQ, respectively, making them effective in ranking diverse retrieval outputs.
-Overall, the results highlight the importance of re-ranking in improving retrieval effectiveness.
-While BM25 provides a strong initial retrieval, the application of neural re-rankers significantly enhances retrieval quality, particularly for Top-1 and Top-5 results.
-The effectiveness of different models varies based on dataset characteristics, with larger transformer-based models excelling in ranking accuracy, while lighter-weight models offer a balance between accuracy and efficiency.
+The sentence-transformer rerankers, particularly GTR-xxl, achieve 42.93% and 39.41% Top-1 accuracy on NQ-Test and WebQ, respectively, making them effective for diverse retrieval tasks.
+4.5. Generator Results
+In this section, we present the performance of Rankify’s generator module when integrated with Retrieval-Augmented Language Modeling (RALM) (Ram et al., 2023).
+We evaluate the effectiveness of various retrieval models, including BM25, Contriever, and DPR, across three open-domain QA datasets: NQ, TriviaQA, and WebQ.
+The experiments utilize multiple LLMs: LLaMA V3 8B, LLaMA V3.1 8B, Gemma 2B/9B, LLaMA 2 13B, and Mistral 7B to assess the robustness of the retrieval and generation pipeline.
+We focus on the Exact Match (EM) metric, which measures the proportion of generated answers that exactly match the ground-truth references.
+Figure 4 provides a consolidated view of the EM scores for all datasets and models.
+From the figure, we observe the following key patterns: For the NQ dataset, DPR consistently outperforms other retrievers, achieving its peak performance with the LLaMA V3 8B model (28.08%).
+On the TriviaQA dataset, BM25 achieves the highest EM score (57.55%) when paired with the Gemma 2 9B model, outperforming both DPR and Contriever.
+For the WebQuestions (WebQ) dataset, DPR again exhibits strong performance, achieving its best EM score (19.83%) with the LLaMA V3 8B model.
 5. Conclusion
@@
 The framework is open-source, actively maintained, and freely available for the research community.
+Future work will focus on incorporating alternative retrieval evaluation metrics such as MAP, Precision, and NDCG, while expanding dataset coverage to include BEIR for broader benchmarking.
+The framework is open-source, actively maintained, and freely available for the research community.